### v2

# Chaos for endomorphisms of completely metrizable groups and linear operators on Fréchet spaces

Abstract.
Using the techniques in topological dynamics, we give a uniform treatment of Li-Yorke chaos, mean Li-Yorke chaos and distributional chaos for continuous endomorphisms of completely metrizable groups, and characterize three kinds of chaos (resp. extreme chaos) in term of the existence of corresponding semi-irregular points (resp. irregular points). We demonstrate some inner automorphisms of Polish groups to illustrate the results. We also apply the results to the chaos theory of continuous linear operators on Fréchet spaces, which improves some results in the literature.
Key words and phrases:
Li-Yorke chaos, mean Li-Yorke chaos, distributional chaos, equicontinuity, sensitivity, endomorphism, completely metrizable group, linear operator, Fréchet space
2020 Mathematics Subject Classification:
Primary: 47A16; Secondary: 37B05
1. Introduction
By a dynamical system, we mean a pair $(X,T)$, where $(X,d)$ is a metric space and $T\colon X\to X$ is continuous. In the seminal paper [LY1975] Li and Yorke initiated the study of chaos theory in dynamical systems. Following the idea in [LY1975], we say that a dynamical system $(X,T)$ is Li-Yorke chaotic if there exists an uncountable subset $K$ of $X$ such that for any two distinct points $x,y\in K$, one has
$\liminf_{n\to\infty}d(T^{n}x,T^{n}y)=0\text{ and }\limsup_{n\to\infty}d(T^{n}x,T^{n}y)>0.$
Li and Yorke proved in [LY1975] that for a continuous interval map $T\colon[0,1]\to[0,1]$ if it has a periodic point of period $3$ then it is Li-Yorke chaotic. Since then, various alternative, but closely related definitions of chaos, have been proposed to capture the complex behaviors of a dynamical system from different perspectives.
To characterize interval maps with positive topological entropy, Schweizer and Smital [SS1994] introduced the notion of distributional chaos as a natural strengthening of Li-Yorke chaos. Later in [SS04] and [BSS05] the authors further introduced three type of distributional chaos. In this paper, we will only focus on distributional chaos of types $1$ and $2$. We say that a dynamical system $(X,T)$ is distributionally chaotic of type $1$ (or of type $2$) if there exists an uncountable subset $K$ of $X$ such that for any two distinct points $x,y\in X$, one has that for every $\varepsilon>0$, the upper density of the set $\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)<\varepsilon\})$ is one and there exists $\delta>0$ such that the upper density of the set $\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})$ is one (or positive).
It is observed by Downarowicz in [D2014] that for dynamical systems on a bounded metric space, distributional chaos of type $2$ is equivalent to another generalization of Li-Yorke chaos, namely mean Li-Yorke chaos, which was coined in [HLY2014]. We say that a dynamical system $(X,T)$ is mean Li-Yorke chaotic if there exists an uncountable subset $K$ of $X$ such that for any two distinct points $x,y\in K$, one has
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0\text{ and }\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)>0.$
For the study of chaos theory in topological dynamics, there are plentiful results about the existence of some kind of chaos and the relationships among various kinds of chaos. We refer the reader to the survey [LY2016] and references therein. Recently, a lot of attention has been paid for chaotic phenomena in linear dynamics, that is in the setting $X$ is a topological vector space (usually Banach or Fréchet space) and $T\colon X\to X$ is a continuous linear operator. We refer the reader to the books [BM2009, GP2011] for a thorough account of linear dynamics.
In a series of papers [BBMP2011, BBMP2015, BBMP2013, BBP2020, BBPW2018], various chaos properties relevant to Li-Yorke chaos were carefully considered, and many very interesting results were found in linear dynamics. One of the important results is that in [BBMP2011], Bermúdez et al. connected Li-Yorke chaotic operators on a Banach space with the existence of irregular points which is discussed in operator theory. Later in [BBMP2015], Bernardes et al. extended this result to Fréchet spaces. They utilized the linear structure and absolutely convex property to construct the uncountable scrambled set. It is natural to ask if the linear structure or locally convex is the key point indeed to produce Li-Yorke chaotic phenomena in linear dynamics.
Every topological vector space is also a topological group. In [C2001], Chan generalized the hypercyclicity criterion in linear dynamics to continuous homomorphisms on a separable completely metrizable topological semigroup. In [M2009], Moothathu studied the weak mixing and mixing properties about a continuous endomorphism of a Hausdorff topological group. In [CW2019] Chan and Walmsley generalized the well-known universality criterion in linear dynamics to the setting of continuous homomorphisms acting on a separable, completely metrizable topological semigroup by constructing a particular universal element. In [BP2022] Burke and Pinheiro obtained necessary and sufficient conditions for a continuous endomorphism on a Polish group to be weakly mixing. In fact, the first chapter of the textbook [B1974] on functional analysis and operator theory is about topological groups. It is natural to investigate the dynamics of continuous endomorphisms of general topological groups. Since both topological dynamics and linear dynamics have plentiful results in chaos theory. So it is interesting and important to connect these two important research areas.
Based on the above reasons, influenced by the seminal paper [HY2002] by Huang and Ye, we take advantage of the techniques in topological dynamics to study the chaos theory in linear dynamics. After distinguishing what is exactly needed in the linear cases, we show that many results hold in the general setting of continuous endomorphisms of completely metrizable groups. We will give a uniform treatment of Li-Yorke chaos, mean Li-Yorke chaos and distributional chaos for continuous endomorphisms of completely metrizable groups in the context. In particular, many results hold for continuous linear operators on F-spaces (completely metrizable topological vector spaces). We will demonstrate some inner automorphisms of the Polish group of all permutations of $\mathbb{Z}$ and the Polish group of all increasing homeomorphisms of $\mathbb{R}$ to illustrate the results. We will apply those results on continuous endomorphisms of completely metrizable groups to continuous linear operators on Fréchet spaces and Banach spaces, which improves some results on the characterizations of chaos via irregular vectors in an abstract and concise way. The paper is organized as follows.
In Section 2 we recall some notations and preliminaries on topological space, topological group, Fréchet space, and dynamical system.
In Section 3, we focus our attention on Li-Yorke chaos. For linear operators on Banach space, Bernardes et al. [BBMP2011] gave a characterize of Li-Yorke chaos by irregular point, which was introduced by Beauzamy in [B1988]. In [BBMP2015] the authors extended this interesting result from Banach space to Fréchet space and systematically studied (dense) Li-Yorke chaos. In this section , we continue to extend this interesting result to a more general situation, that is continuous endomorphisms of completely metrizable groups. To do this, we first study equicontinuity, sensitivity, and Li-Yorke chaos in topological dynamics, and prove the dichotomy of equicontinuity and sensitivity for continuous endomorphisms of completely metrizable groups. Using the group structure and properties of sensitivity, we characterize Li-Yorke chaos (Li-Yorke extreme chaos) in terms of the existence of semi-irregular points (irregular points) for continuous endomorphisms of completely metrizable groups. As applications, we characterize Li-Yorke chaos and Li-Yorke extreme chaos for continuous linear operators on Fréchet spaces.
In Section 4, we focus our attention on mean Li-Yorke chaos. In [BBP2020], Bernardes et al. systematically studied mean (dense) Li-Yorke chaos for linear operators on Banach spaces. In this section, we first study mean equicontinuity, mean sensitivity and mean Li-Yorke chaos in topological dynamics in subsection 4.1. In subsections 4.2 and 4.3, we prove the the dichotomies of mean equicontinuity and mean sensitivity, and give a characterization of (dense) mean Li-Yorke (extreme) chaos for continuous endomorphisms of completely metrizable groups. In particular, utilizing the linear structure and the results in subsections 4.2 and 4.3, we extended the main results in  [BBP2020] concerning (dense) mean Li-Yorke chaos to the Fréchet space setting in subsection 4.4.
In Section 5, we focus our attention on distributional chaos. In [BBMP2013], Bernardes et al. systematically studied (dense) distributionally Li-Yorke chaos of type 1 for linear operators on Fréchet spaces. In [BBPW2018], Bernardes et al. further studied (dense) distributionally Li-Yorke chaos of type 1 and 2 for linear operators on Banach spaces. In this section, we first study mean-L-stability and mean-L-unstability and distributionally Li-Yorke (extreme) chaos of type 1 and 2 in topological dynamics in subsection 5.1. In subsections 5.2 and 5.3, we prove the dichotomies of mean-L-stability and mean-L-unstability, and give a characterization of distributionally Li-Yorke (extreme) chaos of type 1 and 2 for continuous endomorphisms of completely metrizable groups. In particular, Bernardes et al. obtain some sufficient condition for dense distributionally Li-Yorke chaos of type 1 for linear operators in Fréchet spaces and Banach spaces in [BBMP2013]*Theorem 15 and [BBPW2018]*Theorem 33, utilizing the linear structure and the results in subsections 5.2 and 5.3, we give a fuller characterization of dense distributional chaos of type 1 and type 2 for linear operators in Fréchet spaces in subsection 5.4.
2. Preliminaries
Let $\mathbb{N}$ denote the set of positive integers. For a finite subset $A$ of $\mathbb{N}$, $\operatorname{\mathrm{card}}(A)$ denotes the number of elements in $A$. For a subset $A$ of $\mathbb{N}$, the upper density of $A$ is defined by
$\operatorname{\overline{\mathrm{dens}}}(A)=\limsup_{n\to\infty}\frac{\operatorname{\mathrm{card}}(A\cap[1,n])}{n},$
and the lower density of $A$ is defined by
$\operatorname{\underline{\mathrm{dens}}}(A)=\liminf_{n\to\infty}\frac{\operatorname{\mathrm{card}}(A\cap[1,n])}{n}.$
We say that $A$ has density $\operatorname{\mathrm{dens}}(A)$ if $\operatorname{\overline{\mathrm{dens}}}(A)=\operatorname{\underline{\mathrm{dens}}}(A)$, in which case $\operatorname{\mathrm{dens}}(A)$ is equal to this common value.
The following result is a standard exercise in analysis.
Lemma 2.1.
Let $(a_{n})_{n}$ be a sequence in $\mathbb{R}$. Then
  1. (1)
    there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}|a_{n}|=0$ if and only if for every $\varepsilon>0$, $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon|a_{n}|<\varepsilon\})=1$;
  2. (2)
    there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\mathrm{dens}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}|a_{n}|=0$ if and only if for every $\varepsilon>0$, $\operatorname{\mathrm{dens}}(\{n\in\mathbb{N}\colon|a_{n}|<\varepsilon\})=1$;
  3. (3)
    there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}|a_{n}|=\infty$ if and only if for every $M>0$, $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon|a_{n}|>M\})=1$.
2.1. Topological space
Let $X$ be a topological space. A subset $Y$ of $X$ is called a $G_{\delta}$ set if it can be expressed as the countable intersection of open subsets of $X$, a Cantor set if it is homeomorphic to the Cantor ternary set in the unit interval, and a $\sigma$-Cantor set if it can be expressed as the countable union of Cantor subsets of $X$.
We say that a topological space $X$ is completely metrizable if there exists a metric $d$ on $X$ such that $(X,d)$ is a complete metric space and $d$ induces the topology of $X$. The well-known Baire category theorem states that for every completely metrizable space, the intersection of a sequence of open dense sets is still dense. In a completely metrizable space $X$, a subset $Y$ of $X$ is called residual if it contains a dense $G_{\delta}$ subset of $X$.
The following result is due to Mycielski [M1964], which is an important topological tool in this study.
Theorem 2.2 (Mycielski).
Let $X$ be a completely metrizable space without isolated points and $R$ be a dense $G_{\delta}$ subset of $X\times X$. Then for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that
$\bigcup_{j=1}^{\infty}K_{j}\times\bigcup_{j=1}^{\infty}K_{j}\subset R\cup\Delta_{X},$
where $\Delta_{X}=\{(x_{1},x_{2})\in X\times X\colon x_{1}=x_{2}\}$. In addition, if $X$ is separable then we can require that the $\sigma$-Cantor set $\bigcup_{j=1}^{\infty}K_{j}$ is dense in $X$.
2.2. Topological group
Let $(G,\cdot)$ be a group and denote by $e$ the identity of the group $G$. We say that $G$ is a topological group if it is group with a topology such that the group operations are continuous, a (completely) metrizable group if it is a topological group with a (completely) metrizable topology, and a Polish group if it is a separable, completely metrizable group.
By the classical Birkhoff-Kakutani theorem (see e.g. [B1974]*Theorem 6.3), every metrizable topological group $G$ admits a left-invariant compatible metric, that is, a metric $d$ inducing the group’s topology and such that
$d(zx,zy)=d(x,y)$
for all $x,y,z\in G$. But there exist natural examples of completely metrizable groups such that every left-invariant compatible metric is not complete, see e.g. [M2011]. When considering a metrizable topological group $G$, we fix a left-invariant compatible metric $d$ on $G$. Denote by $\operatorname{End}(G)$ the set of all continuous endomorphisms of $G$.
2.3. Topological vector space
Let $X$ be a topological vector space over a field $\mathbb{K}$ (standing for $\mathbb{R}$ or $\mathbb{C}$). Denote by $\mathbf{0}$ the zero vector in $X$. We say that $X$ is an F-space if it is completely metrizable topological vector spaces, and a Fréchet space if it is a locally convex F-space.
Let $X$ be a Fréchet space. Endow $X$ with a separating increasing sequence $(p_{j})_{j}$ of seminorms, which is complete in the metric given by
$d(x,y):=\sum_{j=1}^{\infty}\frac{1}{2^{j}}\min(1,p_{j}(x-y)),\quad x,y\in X.$
As $d$ is a bounded metric on $X$, the boundness with respect to $d$ makes no sense. We need the following boundness with respect to the linear structure of $X$. We say that a subset $Y$ of $X$ is bounded if every neighborhood $V$ of $\mathbf{0}$ in $X$ corresponds a number $s>0$ such that $Y\subset tV$ for every $t>s$. Then it is easy to see that $Y$ is bounded if and only if for any $j\geq 1$, $\sup_{x\in Y}p_{j}(x)<\infty$, and $Y$ is unbounded if and only if there exists $m\in\mathbb{N}$ such that $\sup_{x\in Y}p_{m}(x)=\infty$. Denote by $L(X)$ the set of all continuous linear operators $T\colon X\to X$.
Let $T\in L(X)$. It is clear that with the additive operation of the vector space, $X$ is an abelian completely metrizable group, $d$ is translation-invariant and $T$ is a continuous endomorphism of $X$. So we can apply the results of continuous endomorphisms of completely metrizable groups to continuous linear operators on Fréchet spaces.
2.4. Dynamical system
Recall that a dynamical system is a pair $(X,T)$, where $(X,d)$ is a metric space and $T\colon X\to X$ is continuous. For a point $x\in X$, the orbit of $x$ is the set $\{T^{n}x\colon n\geq 0\}$. A subset $Y$ of $X$ is called $T$-invariant if $TY\subset Y$. Clearly, the orbit of $x$ is $T$-invariant, and the closure of a $T$-invariant set is also $T$-invariant.
If $X$ is a compact metric space, we say that $(X,T)$ is a compact dynamical system. Usually, the study of chaos theory in topological dynamics deals with compact dynamical systems, see the survey [LY2016] and references therein. But in many situations, the results or skills also hold for dynamical systems on completely metrizable spaces. In subsections 3.1, 4.1 and 5.1, we study various versions of equicontinuity, sensitivity and chaos for dynamical systems on completely metrizable spaces. Even though those results must be known for specialist in topological dynamics, we will provide some proofs for completeness.
3. Li-Yorke chaos
3.1. Equicontinuity, sensitivity and Li-Yorke chaos in topological dynamics
Let $(X,T)$ be a dynamical system with a metric $d$ on $X$. A pair $(x,y)\in X\times X$ is called asymptotic if
$\lim_{n\to\infty}d(T^{n}x,T^{n}y)=0,$
and proximal if
$\liminf_{n\to\infty}d(T^{n}x,T^{n}y)=0.$
The asymptotic relation and proximal relation of $(X,T)$, denoted by $\operatorname{Asym}(T)$ and $\operatorname{Prox}(T)$, are the set of all asymptotic pairs and proximal pairs respectively. For $x\in X$, the asymptotic cell of $x$ and proximal cell of $x$ are defined by
$\operatorname{Asym}(T,x):=\{y\in X\colon(x,y)\in\operatorname{Asym}(T)\}$
and
$\operatorname{Prox}(T,x):=\{y\in X\colon(x,y)\in\operatorname{Prox}(T)\},$
respectively.
The $G_{\delta}$ property of the following sets will be critical for later use.
Lemma 3.1.
Let $(X,T)$ be a dynamical system and $\delta>0$. Then
  1. (1)
    for every $x\in X$, proximal cell of $x$ is a $G_{\delta}$ subset of $X$;
  2. (2)
    the proximal relation of $(X,T)$ is a $G_{\delta}$ subset of $X\times X$;
  3. (3)
    for every $x\in X$, the set
    $\Bigl{\{}y\in X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}$
    is a $G_{\delta}$ subset of $X$;
  4. (4)
    the set
    $\Bigl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}$
    is a $G_{\delta}$ subset of $X\times X$.
Proof.
(1) Since
$\operatorname{Prox}(T,x)=\bigcap_{n=1}^{\infty}\Bigl{\{}y\in X\colon\exists k>n\text{ s.t. }d(T^{k}x,T^{k}y)<\tfrac{1}{n}\Bigr{\}},$
by the continuous of $T$, the proximal cell $\operatorname{Prox}(T,x)$ is a $G_{\delta}$ subset of $X$.
(3) Note that
$\Bigl{\{}y\in X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}=\bigcap_{n=1}^{\infty}\Bigl{\{}y\in X\colon\exists k>n\text{ s.t. }d(T^{k}x,T^{k}y)>\delta-\tfrac{1}{n}\Bigr{\}}.$
By the continuous of $T$, the set is a $G_{\delta}$ subset of $X$.
The proof of (2) and (4) are similar with that of (1) and (3). ∎
We say that a dynamical system $(X,T)$ is equicontinuous if for every $\varepsilon>0$, there exists $\delta>0$ such that $d(T^{n}x,T^{n}y)<\varepsilon$ for all $x,y\in X$ with $d(x,y)<\delta$ and $n\in\mathbb{N}$.
Lemma 3.2.
Let $(X,T)$ be a dynamical system. If $(X,T)$ is equicontinuous, then every proximal pair is asymptotic.
Proof.
Let $(x_{0},y_{0})\in X\times X$ be a proximal pair. As $(X,T)$ is equicontinuous, for every $\varepsilon>0$ there exists $\delta>0$ such that for any $x,y\in X$ with $d(x,y)<\delta$, $d(T^{n}x,T^{n}x)<\varepsilon$ for all $n\in\mathbb{N}$. As $(x_{0},y_{0})$ is proximal, there exists $k\in\mathbb{N}$ such that $d(T^{k}x_{0},T^{k}y_{0})<\delta$. Then $d(T^{k+n}x_{0},T^{k+n}y_{0})<\varepsilon$ for all $n\in\mathbb{N}$ and thus $\limsup_{n\to\infty}d(T^{n}x_{0},T^{n}y_{0})\leq\varepsilon$. As $\varepsilon>0$ is arbitrary, $(x_{0},y_{0})$ is asymptotic. ∎
We say that a dynamical system $(X,T)$ is sensitive dependence on initial conditions (sensitive for short) if there exists $\delta>0$ such that for every $x\in X$ and $\varepsilon>0$, there exists $y\in X$ with $d(x,y)<\varepsilon$ and some $n\in\mathbb{N}$ such that $d(T^{n}x,T^{n}y)>\delta$. The number $\delta$ is called a sensitivity constant for $(X,T)$.
Remark 3.3.
Note that if $(X,T)$ is sensitive, then $X$ has no isolated points. If in addition $X$ is completely metrizable, then $X$ is uncountable.
We have the following equivalent conditions of sensitivity.
Lemma 3.4.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is sensitive;
  2. (2)
    there exists $\delta>0$ such that for every $x\in X$, the set
    $\Bigl{\{}y\in X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    there exists $\delta>0$ the set
    $\Bigl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
Proof.
(1$\Rightarrow$(2) Let $\delta_{0}$ be a sensitivity constant of $(X,T)$ and $\delta=\frac{\delta_{0}}{2}$. For any $n\in\mathbb{N}$, let
$X_{n}=\{y\in X\colon\exists k>n\text{ s.t. }d(T^{k}x,T^{k}y)>\delta-\tfrac{1}{n}\}.$
It is clear that $X_{n}$ is open. We first show that $X_{n}$ is dense in $X$. Let $U$ be a nonempty open subset of $X$. By the continuity of $T$, we can shrink $U$ to a nonempty open subset $V$ such that $\operatorname{diam}(T^{i}V)<\delta$ for $i=0,1,\dotsc,n$. Since $(X,T)$ is sensitive, there exist $y_{1},y_{2}\in V$ and $k\in\mathbb{N}$ such that $d(T^{k}y_{1},T^{k}y_{2})>\delta_{0}=2\delta$. By the choice of $V$, we have $k>n$. By the triangle inequality, one has either $d(T^{k}x,T^{k}y_{1})>\delta$ or $d(T^{k}x,T^{k}y_{2})>\delta$. Then either $y_{1}\in X_{n}$ or $y_{2}\in X_{n}$. This shows that $X_{n}\cap U\neq\emptyset$. Then $X_{n}$ is dense in $X$. It is easy to check that
$\Bigl{\{}y\in X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}=\bigcap_{n=1}^{\infty}X_{n},$
which is dense in $X$ by the Baire category theorem.
The implications (2) $\Rightarrow$(3$\Rightarrow$(1) are obvious. ∎
Let $(X,T)$ be a dynamical system. According to the common terms of topological dynamics, we say that a pair $(x,y)\in X\times X$ is Li-Yorke chaotic if
$\liminf_{n\to\infty}d(T^{n}x,T^{n}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}d(T^{n}x,T^{n}y)>0,$
that is, $(x,y)$ is proximal but not asymptotic. A subset $K$ of $X$ is called Li-Yorke scrambled if any two distinct points $x,y\in K$ form a Li-Yorke chaotic pair. We say that the dynamical system $(X,T)$ is Li-Yorke chaotic if there exists an uncountable Li-Yorke scrambled subset of $X$.
Usually, the set of all Li-Yorke chaotic pairs is not a $G_{\delta}$ subset of $X\times X$. Following Lemma 3.1, it is meaningful to consider the set of all Li-Yorke chaotic pairs which separate a given constant.
For a given positive number $\delta$, a pair $(x,y)\in X\times X$ is called Li-Yorke $\delta$-chaotic if
$\liminf_{n\to\infty}d(T^{n}x,T^{n}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta.$
Similarly, we can define Li-Yorke $\delta$-scrambled set and Li-Yorke $\delta$-chaos.
Following [AK2003], we say that a dynamical system $(X,T)$ is Li-Yorke sensitive if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that $(x,y)$ is Li-Yorke $\delta$-chaotic. It is clear that every Li-Yorke sensitive system is sensitive.
Applying the Mycielski theorem, we have the following criterion for the existence of Li-Yorke $\delta$-scrambled sets in the countable union of open subsets. If in addition $X$ is separable, then the Li-Yorke $\delta$-scrambled set can chosen to be dense in $X$.
Proposition 3.5.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke $\delta$-scrambled;
  2. (2)
    the proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is sensitive.
Proof.
(1$\Rightarrow$(2) For any two nonempty open subsets $O_{1}$ and $O_{2}$ of $X$, there exist two Cantor sets $K_{1}\subset O_{1}$ and $K_{2}\subset O_{2}$ such that $K_{1}\cup K_{2}$ is Li-Yorke $\delta$-scrambled. Then $(K_{1}\cup K_{2})\times(K_{1}\cup K_{2})\subset\operatorname{Prox}(T)$ and
$(K_{1}\cup K_{2})\times(K_{1}\cup K_{2})\subset\Bigl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}\cup\Delta_{X}.$
As $O_{1}$ and $O_{2}$ are arbitrary, $\operatorname{Prox}(T)$ is dense in $X\times X$, and by Lemma 3.4 $(X,T)$ is sensitive.
(2$\Rightarrow$(1) As $(X,T)$ is sensitive, by Lemma 3.4, there exists $\delta>0$ such that
$R_{1}=\Bigl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}d(T^{n}x,T^{n}y)\geq\delta\Bigr{\}}.$
is a dense $G_{\delta}$ subset of $X\times X$. Then $R_{1}\cap\operatorname{Prox}(T)$ is also a dense $G_{\delta}$ subset of $X\times X$. By Remark 3.3, $X$ has no isolated points. Applying the Mycielski Theorem to $R_{1}\cap\operatorname{Prox}(T)$, for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that
$\bigcup_{j=1}^{\infty}K_{j}\times\bigcup_{j=1}^{\infty}K_{j}\subset(R_{1}\cap\operatorname{Prox}(T))\cup\Delta_{X}.$
It is clear that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke $\delta$-scrambled. ∎
Motivated by the study of Li-Yorke chaos and distributional chaos for linear operators on Fréchet spaces in [BBMP2013, BBMP2015], we introduce extreme chaos with respect to a continuous pseudometric. From now on in this subsection, we fix a continuous pseudometric $\rho$ on $X$. We say that the dynamical system $(X,T)$ is $\rho$-extremely sensitive if for every $x\in X$ and $\varepsilon>0$, there exists $y\in X$ with $d(x,y)<\varepsilon$ and some $n\in\mathbb{N}$ such that $\rho(T^{n}x,T^{n}y)>\frac{1}{\varepsilon}$.
The idea of the proof of the following result is similar to that of Lemma 3.4 respectively. We leave it to the reader.
Lemma 3.6.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is $\rho$-extremely sensitive;
  2. (2)
    for every $x\in X$, the set
    $\Bigl{\{}y\in X\colon\limsup_{n\to\infty}\rho(T^{n}x,T^{n}y)=\infty\Bigr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    the set
    $\Bigl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}\rho(T^{n}x,T^{n}y)=\infty\Bigr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
Let $(X,T)$ be a dynamical system. A pair $(x,y)\in X\times X$ is called Li-Yorke $\rho$-extremely chaotic if
$\liminf_{n\to\infty}d(T^{n}x,T^{n}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\rho(T^{n}x,T^{n}y)=\infty.$
Similarly, we can define Li-Yorke $\rho$-extremely scrambled set and Li-Yorke $\rho$-extreme chaos and Li-Yorke $\rho$-extreme sensitivity.
The proof of the following result is analogous to that of Proposition 3.5.
Proposition 3.7.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke $\rho$-extremely scrambled;
  2. (2)
    the proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is $\rho$-extremely sensitive.
Remark 3.8.
If the pseudometric $\rho$ is uniformly continuous, then for every $\varepsilon>0$ there exists $\delta>0$ such that for any $x,y\in X$ with $d(x,y)<\delta$, one has $\rho(x,y)<\varepsilon$. So in this case, extreme sensitivity implies sensitivity, and Li-Yorke extreme chaos implies Li-Yorke $\delta$-chaos for some $\delta>0$.
3.2. Li-Yorke chaos for continuous endomorphisms of completely metrizable groups
In [BBMP2011] and [BBMP2015] the authors characterized of Li-Yorke chaos for linear operators on Banach and Fréchet spaces. As explained in the introduction, one of the motivations of this paper is that using the techniques in topological dynamics we show that many results also hold in the general setting of continuous endomorphisms of completely metrizable groups.
In this subsection unless otherwise specified, $G$ denotes a completely metrizable group endowing with a left-invariant compatible metric $d$. Let $T\in\operatorname{End}(G)$, that is, $T$ is a continuous endomorphism of the completely metrizable group $G$. Then $(G,T)$ forms a dynamical system. When stating dynamical properties of $(G,T)$, we will refer the endomorphism $T$ for short.
In [B1988], Beauzamy introduced the concept of irregular vector for continuous linear operators on Banach spaces. In [BBMP2011], Bermúdez et al. characterized Li-Yorke chaos in terms of existence of irregular vectors for continuous linear operators on Banach spaces. Later in [BBMP2015], Bernardes et al. introduced the concept of semi-irregular vector for continuous linear operators on Fréchet spaces. In this paper, we introduce semi-irregular point for continuous endomorphisms of groups. Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called semi-irregular for $T$ if
$\liminf_{n\to\infty}d(T^{n}x,e)=0\qquad\text{and}\qquad\limsup_{n\to\infty}d(T^{n}x,e)>0,$
that is, $(x,e)$ is a Li-Yorke chaotic pair.
By the left-invariance of the metric, we have the following auxiliary observation.
Lemma 3.9.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For every $x\in G$, $\operatorname{Asym}(T,x)=x\cdot\operatorname{Asym}(T,e)$, $\operatorname{Prox}(T,x)=x\cdot\operatorname{Prox}(T,e)$.
  2. (2)
    $\operatorname{Asym}(T,e)$ is dense in $G$ if and only if $\operatorname{Asym}(T)$ is dense in $G\times G$.
  3. (3)
    If $\operatorname{Asym}(T,e)$ is residual in $G$, then $\operatorname{Asym}(T,e)=G$.
  4. (4)
    $\operatorname{Prox}(T,e)$ is dense in $G$ if and only if $\operatorname{Prox}(T)$ is dense in $G\times G$.
  5. (5)
    For every $x,y\in G$, $(x,y)$ is a Li-Yorke chaotic pair if and only if $x^{-1}y$ is a semi-irregular point.
  6. (6)
    If $K$ is a Li-Yorke scrambled subset of $G$, then for any $x\in G$, $xK$ is also a Li-Yorke scrambled set. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of semi-irregular points.
  7. (7)
    For a given $\delta>0$, if $K$ is a Li-Yorke $\delta$-scrambled subset of $G$, then for any $x\in G$, $xK$ is also a Li-Yorke $\delta$-scrambled set.
Proof.
As $d$ is left-invariant, for every $x,y\in G$ and $n\in\mathbb{N}$ one has
$d(T^{n}x,T^{n}y)=d(T^{n}(x^{-1})T^{n}x,T^{n}(x^{-1})T^{n}y)=d(e,T^{n}(x^{-1}y)).$
It is easy to see that (1), (5), (6) and (7) follow from the above formula.
(2) If $\operatorname{Asym}(T,e)$ is dense in $X$, then by (1) for every $x\in G$, $\operatorname{Asym}(T,x)$ is also dense in $X$. As $\operatorname{Asym}(T)=\bigcup_{x\in G}\{x\}\times\operatorname{Asym}(T,x)$, $\operatorname{Asym}(T)$ is dense in $X\times X$.
Now assume that $\operatorname{Asym}(T)$ is dense in $X\times X$. For every nonempty open subset $W$ of $G$, as $G$ is a topological group, there exists an open neighborhood $U$ of $e$ and a nonempty open subset $V$ of $W$ such that $U^{-1}\cdot V\subset W$. Pick a point $(x,y)\in\operatorname{Asym}(T)\cap(U\times V)$. Then $(e,x^{-1}y)\in\operatorname{Asym}(T)$. Therefore, $x^{-1}y\in\operatorname{Asym}(T,e)\cap W$, which implies that $\operatorname{Asym}(T,e)$ is dense in $G$.
(3) For every $x\in G$, $\operatorname{Asym}(T,e)x^{-1}$ is also residual in $X$. Pick $y\in\operatorname{Asym}(T,e)\cap\operatorname{Asym}(T,e)x^{-1}$. Then $(y,e)$ and $(yx,e)$ are asymptotic. As the relation $\operatorname{Asym}(T)$ is transitive, $(y,yx)\in\operatorname{Asym}(T)$ and then $(e,x)\in\operatorname{Asym}(T)$, that is $x\in\operatorname{Asym}(T,e)$.
(4) The proof is similar to (2). ∎
Lemma 3.10.
Let $T\in\operatorname{End}(G)$ and $(n_{k})_{k}$ be an increasing sequence of positive integers. Then
$G_{0}=\Bigl{\{}x\in G\colon\lim_{k\to\infty}T^{n_{k}}x=e\Bigr{\}}$
is a $T$-invariant subgroup of $G$.
Proof.
As $T$ is continuous and $Te=e$, it is easy to verify that $G_{0}$ is $T$-invariant. For every $x,y\in G_{0}$, we have
$\begin{split}\limsup_{k\to\infty}d(T^{n_{k}}(y^{-1}x),e)&=\limsup_{k\to\infty}d(T^{n_{k}}x,T^{n_{k}}y)\\ &\leq\lim_{k\to\infty}d(T^{n_{k}}x,e)+\lim_{k\to\infty}d(e,T^{n_{k}}y)\\ &=0.\end{split}$
Then $G_{0}$ is a subgroup of $G$. ∎
In [AY1980], Auslander and Yorke proved that for a compact dynamical system if it is minimal (that is, it does not contain any proper closed invariant subsets) then it is either equicontinuous or sensitive. We have the following dichotomy of equicontinuity and sensitivity for continuous endomorphisms of completely metrizable groups.
Theorem 3.11.
Let $T\in\operatorname{End}(G)$. Then $T$ is either equicontinuous or sensitive.
Proof.
Let $e$ be the identity of $G$. We have the following two cases:
Case 1: for any $\varepsilon>0$, there exists $\delta>0$ such that for every $x\in G$ with $d(x,e)<\delta$, we have $d(T^{n}x,e)<\varepsilon$ for all $n\in\mathbb{N}$. Let $x_{1},x_{2}\in G$. If $d(x_{1},x_{2})<\delta$, by the left-invariance of $d$ we have $d(x_{1},x_{2})=d(x_{2}^{-1}x_{1},e)<\delta$. Then for every $n\in\mathbb{N}$,
$d(T^{n}x_{1},T^{n}x_{2})=d(T^{n}(x_{2}^{-1})T^{n}x_{1},e)=d(T^{n}(x_{2}^{-1}x_{1}),e)<\varepsilon.$
This implies that $T$ is equicontinuous.
Case 2: there exists $\delta>0$ such that for every $\varepsilon>0$ there exists $y_{\varepsilon}\in G$ with $d(y_{\varepsilon},e)<\varepsilon$ and $n\in\mathbb{N}$ such that $d(T^{n}y_{\varepsilon},e)>\delta$. For every $x\in G$ and $\varepsilon>0$, we have $d(xy_{\varepsilon},x)<\varepsilon$ and
$d(T^{n}(xy_{\varepsilon}),T^{n}x)=d(T^{n}(x^{-1})T^{n}(xy_{\varepsilon}),e)=d(T^{n}y_{\varepsilon},e)>\delta.$
Thus $T$ is sensitive. ∎
Combining Lemma 3.2 and Theorem 3.11, we have the following corollary.
Corollary 3.12.
Let $T\in\operatorname{End}(G)$. If there exists a Li-Yorke chaotic pair, then $T$ is sensitive.
Now we have the following characterization about the existence of a dense set of semi-irregular points for continuous endomorphisms of completely metrizable groups.
Theorem 3.13.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of semi-irregular points;
  2. (2)
    $T$ admits a residual set of semi-irregular points;
  3. (3)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke $\delta$-scrambled;
  4. (4)
    the proximal relation of $T$ is dense in $G\times G$ and $T$ is sensitive;
  5. (5)
    the proximal cell of $e$ is dense in $G$ and there exists $x\in G$ such that
    $\limsup_{n\to\infty}d(T^{n}x,e)>0.$
Proof.
The implications (2) $\Rightarrow$(1$\Rightarrow$(5) are clear.
(5$\Rightarrow$(4) By Lemma 3.1 (1), $\operatorname{Prox}(T,e)$ is residual in $X$. If $\operatorname{Prox}(T,e)=\operatorname{Asym}(T,e)$, then by Lemma 3.9 (3), one has $\operatorname{Asym}(T,e)=G$, which is a contradiction. So $\operatorname{Prox}(T,e)\setminus\operatorname{Asym}(T,e)\neq\emptyset$, that is $\operatorname{Prox}(T,e)$ contains a semi-irregular point. Now by Corollary 3.12, $T$ is sensitive.
(3) $\Leftrightarrow$ (4) It follows from Proposition 3.5.
(4) $\Rightarrow$ (2) As $T$ is sensitive, by Lemma 3.4 there exists $\delta>0$ such that
$\Bigl{\{}x\in G\colon\limsup_{n\to\infty}d(T^{n}x,e)\geq\delta\Bigr{\}}$
is a dense $G_{\delta}$ subset of $G$. Let
$G_{0}=\operatorname{Prox}(T,e)\cap\Bigl{\{}x\in G\colon\limsup_{n\to\infty}d(T^{n}x,e)\geq\delta\Bigr{\}}.$
Then $G_{0}$ is a dense $G_{\delta}$ subset of $G$ consisting of semi-irregular points. ∎
According to Theorem 3.13 (5), we have the following consequences.
Corollary 3.14.
Let $T\in\operatorname{End}(G)$. If the proximal cell of $e$ is dense in $G$ and $T$ has a non-trivial periodic point, then it has a dense set of semi-irregular points.
Applying Theorem 3.13, we have the following characterization of Li-Yorke chaos for continuous endomorphisms of completely metrizable groups.
Theorem 3.15.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a semi-irregular point;
  2. (2)
    $T$ admits a Li-Yorke chaotic pair;
  3. (3)
    $T$ is Li-Yorke chaotic;
  4. (4)
    $T$ is Li-Yorke sensitive;
  5. (5)
    the restriction of $T$ to some invariant closed subgroup $\widetilde{G}$ has a dense set of semi-irregular points.
Proof.
The implications (3$\Rightarrow$(2$\Rightarrow$(1) and (4$\Rightarrow$(2) are clear.
(5$\Rightarrow$(3) It follows from Theorem 3.13.
(1$\Rightarrow$(5) Let $x$ be a semi-irregular point. Then there exists an increasing sequence $(n_{k})_{k}$ of positive integers such that $\lim_{k\to\infty}T^{n_{k}}x=e$. Let
$G_{0}=\Bigl{\{}y\in G\colon\lim_{k\to\infty}T^{n_{k}}y=e\Bigr{\}}.$
It is clear that $x\in G_{0}$. By Lemma 3.10 $G_{0}$ is a $T$-invariant subgroup of $G$. Let $\widetilde{G}$ be the closure of $G_{0}$. Then $\widetilde{G}$ is also a $T$-invariant subgroup of $G$. As $(x,e)$ is Li-Yorke chaotic for $T|_{\widetilde{G}}$, by Corollary 3.12 $T|_{\widetilde{G}}$ is sensitive. It is clear that $G_{0}\subset\operatorname{Prox}(T|_{\widetilde{G}},e)$. So $\operatorname{Prox}(T|_{\widetilde{G}})$ is dense in $\widetilde{G}\times\widetilde{G}$. Now (5) follows from Theorem 3.13.
(5$\Rightarrow$(4) Applying Theorem 3.13 to $T|_{\widetilde{G}}$, there exists $\delta>0$ such that
$G_{0}=\operatorname{Prox}(T,e)\cap\Bigl{\{}x\in\widetilde{G}\colon\limsup_{n\to\infty}d(T^{n}x,e)\geq\delta\Bigr{\}}.$
is residual in $\widetilde{G}$. For every $x\in G$ and $y\in xG_{0}$, $(x,y)$ is Li-Yorke $\delta$-chaotic. Note that $x$ is an accumulation point of $xG_{0}$. Then $T$ is Li-Yorke sensitive. ∎
Remark 3.16.
(a) For Theorem 3.13, if in addition $G$ is separable, then the assertion (3) can be replace by the following assertion:
  1. (${}^{\prime}$)
    there exists a dense $\sigma$-Cantor Li-Yorke $\delta$-scrambled subset of $G$ for some $\delta>0$.
(b) In the proof of (1$\Rightarrow$(5) in Theorem 3.15, if we let $G_{1}$ be the countable group generated by the orbit of $x$, then $G_{1}$ is a $T$-invariant subgroup of $G_{0}$. Let $\widetilde{G_{1}}$ be the closure of $G_{1}$. Then $\widetilde{G_{1}}$ is a $T$-invariant closed separable subgroup of $G$. So the assertion (5) in Theorem 3.15 can be replace by the following assertion:
  1. (${}^{\prime}$)
    the restriction of $T$ to some invariant closed separable subgroup $\widetilde{G}$ has a dense set of semi-irregular points.
Due to the group structure, the following result shows that the existence of a residual Li-Yorke scrambled set implies that the whole space is a Li-Yorke scrambled set.
Proposition 3.17.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a residual Li-Yorke scrambled set;
  2. (2)
    every point in $G\setminus\{e\}$ is semi-irregular for $T$;
  3. (3)
    $X$ is a Li-Yorke scrambled set for $T$.
Proof.
The implications (2) $\Rightarrow$ (3) $\Rightarrow$(1) are clear.
(1) $\Rightarrow$ (2) Let $K$ be a residual Li-Yorke scrambled subset of $X$. For every $x\in G\setminus\{e\}$, $K\cap Kx^{-1}$ is also residual. Pick $y\in K\cap Kx^{-1}$. One has $(y,yx)$ is Li-Yorke chaotic. Then $(e,x)$ is also Li-Yorke chaotic, which shows that $x$ is a semi-irregular point. ∎
Remark 3.18.
It is worth to notice that a point $x\in G$ is semi-irregular for a continuous endomorphism of $T$ if and only if the orbit of $x$ has a subsequence which is convergent to the identity $e$ but the whole orbit is not convergent. So being semi-irregular is a topological property and does not dependent on the compatible metric. Even though in the beginning of this subsection we fix a left-invariant compatible metric $d$ on $G$, Li-Yorke chaos in Theorems 3.13 and 3.15 in deed happens for any left-invariant compatible metric on $G$.
Now we present two kinds of inner automorphisms of Polish groups which are densely Li-Yorke chaotic.
Example 3.19.
Let $\mathbb{S}(\mathbb{Z})$ be the group of all permutations of the integers $\mathbb{Z}$. With respect to the topology of pointwise convergence $\mathbb{S}(\mathbb{Z})$ is a Polish topological group. Let $\sigma\in\mathbb{S}(\mathbb{Z})$ be the unit shift permutation, that is $\sigma(n)=n-1$ for each $n\in\mathbb{Z}$. Define $T\colon\mathbb{S}(\mathbb{Z})\to\mathbb{S}(\mathbb{Z})$ by $T(f)=\sigma f\sigma^{-1}$. Then for every $f\in\mathbb{S}(\mathbb{Z})$, $n,k\in\mathbb{Z}$, $T^{n}(f)(k)=f(k+n)-n$. It is shown in [M2009]*Example 1 that $(\mathbb{S}(\mathbb{Z}),T)$ is mixing.
Let $A$ be the subgroup of $\mathbb{S}(\mathbb{Z})$ consisting of the permutations which fix all but a finite set in $\mathbb{Z}$. It is easy to see that $A$ is dense in $\mathbb{S}(\mathbb{Z})$ and for any $f\in A$, $T^{n}(f)$ converges in $\mathbb{S}(\mathbb{Z})$ to the identity permutation as $n\to\infty$. Then $A\times A$ is contained in the asymptotic relation with respect to any compatible metric on $\mathbb{S}(\mathbb{Z})$. It is clear that $T(\sigma)=\sigma$. Then by Theorem 3.13 and Remark 3.16, $(\mathbb{S}(\mathbb{Z}),T)$ is densely Li-Yorke chaotic for any left-invariant compatible metric on $\mathbb{S}(\mathbb{Z})$.
An precise left-invariant compatible metric $d$ on $\mathbb{S}(\mathbb{Z})$ is defined by $d(f,g)=\frac{1}{2^{k}}$ where $k=\min\{|n|\colon f(n)\neq g(n),n\in\mathbb{Z}\}$ for any $f\neq g\in\mathbb{S}(\mathbb{Z})$. Define $\rho\colon\mathbb{S}(\mathbb{Z})\times\mathbb{S}(\mathbb{Z})\to\mathbb{R}$ by $\rho(f,g)=|f(0)-g(0)|$. Then $\rho$ is a continuous pseudometric on $\mathbb{S}(\mathbb{Z})$. It is easy to see that $(\mathbb{S}(\mathbb{Z}),T)$ is $\rho$-extremely sensitive. By Proposition 3.7, $(\mathbb{S}(\mathbb{Z}),T)$ is densely Li-Yorke $\rho$-extremely chaotic.
Example 3.20.
Let $H_{+}(\mathbb{R})$ be the group of all increasing homeomorphisms of $\mathbb{R}$. With respect to the compact-open topology $H_{+}(\mathbb{R})$ is a Polish topological group. Let $\sigma\in H_{+}(\mathbb{R})$ be the unit translation, that is $\sigma(x)=x-1$ for each $x\in\mathbb{R}$. Define $T\colon H_{+}(\mathbb{R})\to H_{+}(\mathbb{R})$ by $T(f)=\sigma f\sigma^{-1}$. Then for every $f\in H_{+}(\mathbb{R})$, $n\in\mathbb{Z}$ and $x\in\mathbb{R}$, $T^{n}(f)(x)=f(x+n)-n$. It is shown in [M2009]*Example 3 that $(H_{+}(\mathbb{R}),T)$ is mixing.
Let $A$ be the subgroup of $H_{+}(\mathbb{R})$ consisting of the increasing homeomorphisms which fix all but a bounded set in $\mathbb{R}$. It is easy to see that $A$ is dense in $H_{+}(\mathbb{R})$ and for any $f\in A$, $T^{n}(f)$ converges in $H_{+}(\mathbb{R})$ to the identity as $n\to\infty$. Then $A\times A$ is contained in the asymptotic relation with respect to any compatible metric on $H_{+}(\mathbb{R})$. It is clear that $T(\sigma)=\sigma$. Then by Theorem 3.13 and Remark 3.16, $(H_{+}(\mathbb{R}),T)$ is densely Li-Yorke chaotic for any left-invariant compatible metric on $H_{+}(\mathbb{R})$.
An precise left-invariant compatible metric $d$ on $H_{+}(\mathbb{R})$ is defined by $d(f,g)=2^{-a}$, where $a=\inf\{|x|\colon f(x)\neq g(x),x\in\mathbb{R}\}$ for any $f\neq g\in H_{+}(\mathbb{R})$. Define $\rho\colon H_{+}(\mathbb{R})\times H_{+}(\mathbb{R})\to\mathbb{R}$ by $\rho(f,g)=|f(0)-g(0)|$. Then $\rho$ is a continuous pseudometric on $H_{+}(\mathbb{R})$. It is easy to see that $(H_{+}(\mathbb{R}),T)$ is $\rho$-extremely sensitive. By Proposition 3.7, $(H_{+}(\mathbb{R}),T)$ is densely Li-Yorke $\rho$-extremely chaotic.
3.3. Li-Yorke extreme chaos for continuous endomorphisms of completely metrizable groups
With the same as in the previous subsection, $G$ denotes a completely metrizable group endowing with a left-invariant compatible metric $d$. As every Fréchet space has a separating increasing sequence of seminorms, we consider the counterpart, pseudometric, for metrizable groups. In this subsection we also fix a left-invariant continuous pseudometric $\rho$ on $G$.
Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called irregular for $T$ if
$\liminf_{n\to\infty}d(T^{n}x,e)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\rho(T^{n}x,e)=\infty,$
that is, $(x,e)$ is a Li-Yorke extremely chaotic pair.
Similar to Lemma 3.9, we have the following observation.
Lemma 3.21.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For every $x,y\in G$, $(x,y)$ is a Li-Yorke extremely chaotic pair if and only if $x^{-1}y$ is an irregular point.
  2. (2)
    If $K$ is a Li-Yorke extremely scrambled subset of $G$, then for any $x\in G$, $xK$ is also a Li-Yorke extremely scrambled set. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of irregular points.
Proposition 3.22.
Let $T\in\operatorname{End}(G)$. Then $T$ is extremely sensitive if and only if there exists a sequence $(y_{k})_{k}$ in $X$ such that $\lim_{k\to\infty}y_{k}=e$ and $\limsup_{k\to\infty}\rho(T^{k}y_{k},e)=\infty$.
Proof.
The necessity follows from the definition immediately. Now we prove the sufficiency. For any $\varepsilon>0$ and $x\in G$, pick $k\in\mathbb{N}$ such that $d(y_{k},e)<\varepsilon$ and $\rho(T^{k}y_{k},e)>\frac{1}{\varepsilon}$. By left-invariance of $d$, we have $d(xy_{k},x)=d(y_{k},e)<\varepsilon$ and $\rho(T^{k}(xy_{k}),T^{k}x)=\rho(T^{k}y_{k},e)>\frac{1}{\varepsilon}$. Then $T$ is extremely sensitive. ∎
Proposition 3.23.
Let $T\in\operatorname{End}(G)$. If there exists an irregular point, then $T$ is extremely sensitive.
Proof.
Let $y\in G$ be an irregular point. Then there exists a sequence $(i_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}T^{i_{k}}y=e$ and $\limsup_{n\to\infty}\rho(T^{n}y,e)=\infty$. Note that for any $k\in\mathbb{N}$, $\limsup_{n\to\infty}\rho(T^{n}(T^{i_{k}}y),e)=\infty$. Then by Proposition 3.22, $T$ is extremely sensitive. ∎
Similar to Theorem 3.13, we have the following characterization about the existence of a dense set of irregular points for continuous endomorphisms of completely metrizable groups.
Theorem 3.24.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of irregular points;
  2. (2)
    $T$ admits a residual set of irregular points;
  3. (3)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke extremely scrambled;
  4. (4)
    the proximal relation of $T$ is dense in $G\times G$ and $T$ is extremely sensitive.
Proof.
The implication (2$\Rightarrow$(1) is clear.
(1$\Rightarrow$(4) Let $A$ be a dense set of irregular points. Then $A\subset\operatorname{Prox}(T,e)$. By Lemma 3.9, $\operatorname{Prox}(T)$ is dense in $G\times G$. Lemma 3.23 ensures that $T$ is extremely sensitive.
(3$\Leftrightarrow$(4) It follows from Proposition 3.7.
(4$\Rightarrow$(2) The set of irregular points equals to the intersection
$\operatorname{Prox}(T,e)\cap\Bigl{\{}x\in X\colon\limsup_{n\to\infty}\rho(T^{n}x,e)=\infty\Bigr{\}}$
By Lemma 3.9, $\operatorname{Prox}(T,e)$ is a dense $G_{\delta}$ subset of $G$. As $T$ is extremely sensitive, by Lemma 3.6, the set
$\Bigl{\{}x\in X\colon\limsup_{n\to\infty}\rho(T^{n}x,e)=\infty\Bigr{\}}$
is also a dense $G_{\delta}$ subset of $G$. Then the set of irregular points is residual. ∎
Using the same arguments as in the proof of Theorem 3.15, applying Theorem 3.24 we have the following characterization of Li-Yorke extremely chaos for continuous endomorphisms of completely metrizable groups.
Theorem 3.25.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits an irregular point;
  2. (2)
    $T$ admits a Li-Yorke extremely chaotic pair;
  3. (3)
    $T$ is Li-Yorke extremely chaotic;
  4. (4)
    $T$ is Li-Yorke extremely sensitive;
  5. (5)
    the restriction of $T$ to some invariant closed subgroup $\widetilde{G}$ has a dense set of irregular points.
Remark 3.26.
(a) For Theorem 3.24, if in addition $G$ is separable, then the assertion (3) can be replace by the following assertion:
  1. (${}^{\prime}$)
    there exists a dense $\sigma$-Cantor Li-Yorke extremely scrambled subset of $G$.
(b) In the proof of (1$\Rightarrow$(5) in Theorem 3.25, if we let $G_{1}$ be the countable group generated by the orbit of $x$, then $G_{1}$ is a $T$-invariant subgroup of $G_{0}$. Let $\widetilde{G_{1}}$ be the closure of $G_{1}$. Then $\widetilde{G_{1}}$ is a $T$-invariant closed separable subgroup of $G$. So the assertion (5) in Theorem 3.25 can be replace by the following assertion:
  1. (${}^{\prime}$)
    the restriction of $T$ to some invariant closed separable subgroup $\widetilde{G}$ has a dense set of irregular points.
3.4. Li-Yorke chaos for continuous linear operators on Fréchet spaces
In this subsection unless otherwise specified, $X$ denotes an arbitrary infinite-dimensional Fréchet space. Let $(p_{j})_{j}$ be a separating increasing sequence of seminorms on $X$. The metric $d$ on $X$ is given by
$d(x,y):=\sum_{j=1}^{\infty}\frac{1}{2^{j}}\min(1,p_{j}(x-y)),\quad x,y\in X.$
Though Li-Yorke chaos for linear operators on Banach and Fréchet spaces has been characterized in [BBMP2011] and [BBMP2015], in this subsection using results in the subsections 3.2 and 3.3 we will reprove some results in [BBMP2011] and [BBMP2015] using different approaches, and also slightly generalize some of them.
First we have the following auxiliary result.
Lemma 3.27.
Let $T\in L(X)$ and $(n_{k})_{k}$ be an increasing sequence of positive integers. Then
$X_{0}:=\Bigl{\{}x\in X\colon\lim_{k\to\infty}T^{n_{k}}x=\mathbf{0}\Bigr{\}}$
is a $T$-invariant subspace of $X$.
Proof.
As $T$ is continuous and $T\mathbf{0}=\mathbf{0}$, it is easy to verify that $X_{0}$ is $T$-invariant. For every $x,y\in X_{0}$ and $\alpha,\beta\in\mathbb{K}$, we have
$\begin{split}\lim_{k\to\infty}T^{n_{k}}(\alpha x+\beta y)&=\alpha\lim_{k\to\infty}T^{n_{k}}x+\beta\lim_{k\to\infty}T^{n_{k}}y\\ &=\mathbf{0}.\end{split}$
Then $X_{0}$ is a subspace of $X$. ∎
Recall that $T\in L(X)$ is equicontinuous as a dynamical property if and only if the family $\{T^{n}\colon n\in\mathbb{N}\}$ of operators is equicontinuous in the theory of functional analysis. Then the following result follows from the well-known Banach-Steinhaus Theorem directly.
Proposition 3.28.
Let $T\in L(X)$. Then $T$ is equicontinuous if and only if for every $x\in X$, the orbit of $x$ is bounded.
Remark 3.29.
Assume in addition that $X$ is Banach space. Then $T\in L(X)$ is equicontinuous if and only if $T$ is power-bounded, that is $\sup_{n}\|T^{n}\|<\infty$.
Each seminorm $p_{j}$ induces a continuous left-invariant pseudometric $d_{j}$ defined by $d_{j}(x,y)=p_{j}(x-y)$. Let $T\in L(X)$ and $m\in\mathbb{N}$. A vector $x\in X$ is called $m$-irregular if it is irregular with respect to the seminorm $p_{m}$. A pair $(x,y)\in X\times X$ is called Li-Yorke $m$-extremely chaotic if it is Li-Yorke extremely chaotic with respect to the seminorm $p_{m}$. Similarly, we can define $m$-extreme sensitivity, Li-Yorke $m$-extremely scrambled set, Li-Yorke $m$-extreme chaos and Li-Yorke $m$-extreme sensitivity.
Remark 3.30.
[…]
$\liminf_{n\to\infty}\lVert T^{n}x\rVert=0\qquad\text{and}\qquad\limsup_{n\to\infty}\lVert T^{n}x\rVert=\infty,$
a pair $(x,y)\in X\times X$ is Li-Yorke extremely chaotic if and only if
$\liminf_{n\to\infty}\lVert T^{n}x-T^{n}y\rVert=0\qquad\text{and}\qquad\limsup_{n\to\infty}\lVert T^{n}x-T^{n}y\rVert=\infty.$
Proposition 3.31.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is sensitive;
  2. (2)
    there exists an unbounded orbit;
  3. (3)
    there exists a bounded sequence $(y_{k})_{k}$ in $X$ such that $(T^{k}y_{k})_{k}$ is unbounded;
  4. (4)
    $T$ is $m$-extremely sensitive for some $m\in\mathbb{N}$.
Proof.
(1$\Leftrightarrow$(2) It follows immediately from Theorem 3.11 and Proposition 3.28.
The implications (2$\Rightarrow$(3) and $(4)\Rightarrow$(1) are clear.
(3$\Rightarrow$(4) As $(T^{k}y_{k})_{k}$ is unbounded, there exists $m\in\mathbb{N}$ such that $\sup_{k}p_{m}(T^{k}y_{k})=\infty$. For each $n>0$, there exists $N\in\mathbb{N}$ such that $d(\frac{1}{N}y_{k},\mathbf{0})<\frac{1}{n}$ for all $k\in\mathbb{N}$. Pick $k\in\mathbb{N}$ such that $p_{m}(T^{k}y_{k})>nN$. Then $p_{m}(T^{k}(\frac{1}{N}y_{k}))>n$. By Proposition 3.22, $T$ is $m$-extremely sensitive. ∎
Combining Proposition 3.31, Theorems 3.13 and 3.24, we have the following characterization about the existence of a dense set of semi-irregular vectors for continuous linear operators on Fréchet spaces, which generalizes [BBMP2015]*Theorem 10 from separable Fréchet spaces to general Fréchet spaces.
Theorem 3.32.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of semi-irregular vectors;
  2. (2)
    $T$ admits a residual set of $m$-irregular vectors for some $m\in\mathbb{N}$;
  3. (3)
    there exists $m\in\mathbb{N}$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is Li-Yorke $m$-extremely scrambled;
  4. (4)
    the proximal relation of $T$ is dense in $X\times X$ and $T$ is sensitive;
  5. (5)
    the proximal cell of $\mathbf{0}$ is dense in $X$ and there exists $x\in X$ such that
    $\limsup_{n\to\infty}d(T^{n}x,\mathbf{0})>0.$
Applying Theorem 3.32, we obtain the following characterization of Li-Yorke $m$-extreme chaos for continuous linear operators on Fréchet spaces, which slightly strengthens [BBMP2015]*Theorem 9 and the equivalent assertion (5) seems new.
Theorem 3.33.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a semi-irregular vector;
  2. (2)
    $T$ admits a Li-Yorke chaotic pair;
  3. (3)
    $T$ admits an $m$-irregular vector for some $m\in\mathbb{N}$;
  4. (4)
    $T$ is Li-Yorke $m$-extremely chaotic for some $m\in\mathbb{N}$;
  5. (5)
    $T$ is Li-Yorke $m$-extremely sensitive for some $m\in\mathbb{N}$;
  6. (6)
    the restriction of $T$ to some closed $T$-invariant subspace $\widetilde{X}$ has a residual set of $m$-irregular points for some $m\in\mathbb{N}$.
Proof.
The implications (4) $\Rightarrow$(3$\Rightarrow$(2$\Rightarrow$(1) and (5) $\Rightarrow$(1) are clear.
(6$\Rightarrow$(4) It follows from Theorem 3.32.
(1$\Rightarrow$(6) Let $x$ be a semi-irregular vector. There exists an increasing sequence of positive integers $(n_{k})$ such that $\lim_{k\to\infty}T^{n_{k}}x=\mathbf{0}$. Let
$X_{0}=\Bigl{\{}y\in X\colon\lim_{k\to\infty}T^{n_{k}}y=\mathbf{0}\Bigr{\}}.$
It is clear that $x\in X_{0}$. By Lemma 3.27 $X_{0}$ is a $T$-invariant subspace of $X$. Let $\widetilde{X}$ be the closure of $X_{0}$. Then $\widetilde{X}$ is also a $T$-invariant subspace of $X$. As $(x,e)$ is Li-Yorke chaotic for $T|_{\widetilde{X}}$, by Corollary 3.12 $T|_{\widetilde{X}}$ is sensitive. It is clear that $X_{0}\subset\operatorname{Prox}(T|_{\widetilde{X}},e)$. So $\operatorname{Prox}(T|_{\widetilde{X}})$ is dense in $\widetilde{X}\times\widetilde{X}$. Now (6) follows from Theorem 3.32.
(6$\Rightarrow$(5) Applying Theorem 3.32 to $T|_{\widetilde{X}}$, there exists a residual subset $X_{0}$ of $\widetilde{X}$ consisting of $m$-irregular vector for some $m\in\mathbb{N}$. For every $x\in X$ and $y\in x+X_{0}$, $(x,y)$ is Li-Yorke $m$-extremely chaotic. Note that $x$ is an accumulation point of $x+X_{0}$. Then $T$ is Li-Yorke $m$-extremely sensitive. ∎
By the proof of (1$\Rightarrow$(6) in Theorem 3.33, we have the following consequence.
Corollary 3.34.
Let $T\in L(X)$. Then the set of $m$-irregular vectors for some $m\in\mathbb{N}$ is dense in the set of semi-irregular vectors.
Following [BBMP2015], we say that $T\in L(X)$ satisfies the Li-Yorke chaos criterion if there exists a subset $X_{0}$ of $X$ with the following properties:
  1. (1)
    for every $x\in X_{0}$, $(T^{n}x)_{n}$ has a subsequence converging to $\mathbf{0}$;
  2. (2)
    there is a bounded sequence $(a_{n})_{n}$ in $\overline{\operatorname{span}(X_{0})}$ such that the sequence $(T^{n}a_{n})_{n}$ is unbounded.
Using the ideas developed in this subsection, we can give a new proof of the following result.
Proposition 3.35 ([BBMP2015]*Theorem 15).
Let $T\in L(X)$. Then $T$ is Li-Yorke chaotic if and only if $T$ satisfies the Li-Yorke chaos criterion.
Proof.
If $T$ is Li-Yorke chaotic, by Theorem 3.33 the restriction of $T$ to some invariant closed $T$-invariant subspace $\widetilde{X}$ has a residual set of $m$-irregular points for some $m\in\mathbb{N}$. Let $X_{0}=\operatorname{Prox}(T_{\widetilde{X}},\mathbf{0})$. Then by Theorem 3.32 and Proposition 3.31, $X_{0}$ is as required for the Li-Yorke chaos criterion.
Now assume that $T$ satisfies the Li-Yorke chaos criterion. If $X_{0}\setminus\operatorname{Asym}(T,e)\neq\emptyset$, that is there exists a semi-irregular vector in $X_{0}$, by Theorem 3.33 $T$ is Li-Yorke chaotic. Otherwise, $X_{0}\subset\operatorname{Asym}(T,e)$. In this case let $\widetilde{X}$ be the closure of $\operatorname{Asym}(T,e)$. By Lemma 3.27 $\widetilde{X}$ is a $T$-invariant closed subspace of $X$. Then the proximal cell of $e$ with respect to $T|_{\widetilde{X}}$ is dense in $\widetilde{X}$. By Proposition 3.31, $T|_{\widetilde{X}}$ is sensitive. Applying Theorem 3.32 to $T|_{\widetilde{X}}$, one has that $T$ is Li-Yorke chaotic. ∎
Let $T\in L(X)$. A vector subspace $Y$ of $X$ is called an irregular manifold for $T$ if every vector $y\in Y\setminus\{0\}$ is irregular for $T$. The following result is inspired by [BBMP2015]*Theorem 20 which states that if the asymptotic cell of $\mathbf{0}$ is dense in $X$ and $T$ is sensitive then $T$ admits a dense irregular manifold. The proof here is different from that of [BBMP2015]*Theorem 20.
Theorem 3.36.
Let $X$ be a separable Fréchet space and $T\in L(X)$. If the asymptotic cell of $\mathbf{0}$ is dense in $X$, then either the asymptotic cell of $\mathbf{0}$ is $X$ or $T$ admits a dense irregular manifold.
Proof.
Note that $\operatorname{Prox}(T,\mathbf{0})$ is residual in $X$, as $\operatorname{Asym}(T,\mathbf{0})\subset\operatorname{Prox}(T,\mathbf{0})$. If $\operatorname{Asym}(T,\mathbf{0})=\operatorname{Prox}(T,\mathbf{0})$, then by Lemma 3.9 (3), $\operatorname{Asym}(T,\mathbf{0})=X$. Otherwise $\operatorname{Prox}(T,\mathbf{0})\setminus\operatorname{Asym}(T,\mathbf{0})\neq\emptyset$, that is $T$ admits a semi-irregular vector. Then by Corollary 3.12 $T$ is sensitive. Combining Lemma 3.6 and Proposition 3.31, there exists $m\in\mathbb{N}$ such that the set
$D:=\Bigl{\{}x\in X\colon\limsup_{n\to\infty}p_{m}(T^{n}x)=\infty\Bigr{\}}$
is a dense $G_{\delta}$ subset of $X$. We need the following claim.
Claim: For any increasing sequence $(s_{k})_{k}$ in $\mathbb{N}$, the set
$P(s_{k})=\Bigl{\{}x\in X\colon\liminf_{k\to\infty}d(T^{s_{k}}x,\mathbf{0})=0\Bigr{\}}$
is a dense $G_{\delta}$ subset of $X$.
Proof of the Claim.
As $\operatorname{Asym}(T,\mathbf{0})\subset P(s_{k})$, $P(s_{k})$ is dense in $X$. Note that
$P(s_{k})=\bigcap_{n=1}^{\infty}\bigl{\{}x\in X\colon\exists k>n\text{ s.t. }d(T^{s_{k}}x,\mathbf{0})<\tfrac{1}{n}\bigr{\}}.$
Then it is easy to see that $P(s_{k})$ is a $G_{\delta}$ subset of $X$. ∎
Fix a dense sequence $(y_{i})_{i}$ in $X$. We will construct inductively a sequence $(x_{i})_{i}$ in $X$ such that $d(x_{i},y_{i})<\frac{1}{i}$ and $\operatorname{span}\{x_{i}\colon i\in\mathbb{N}\}$ is an irregular manifold.
By the completeness of $X$, $X_{1}:=P(k)\cap D$ is a dense $G_{\delta}$ subset of $X$. We pick $x_{1}\in X_{1}$ with $d(x_{1},y_{1})<1$. Then there exist two increasing sequences $(s^{(1,1)}_{k})_{k}$ and $(t^{(1)}_{k})_{k}$ in $\mathbb{N}$ such that
$\lim_{k\to\infty}d(T^{s^{(1,1)}_{k}}x_{1},\mathbf{0})=0\text{ and }\lim_{k\to\infty}p_{m}(T^{t^{(1)}_{k}}x_{1})=\infty.$
It is clear that $\mathbb{K}x_{1}$ is an irregular manifold.
Assume that $x_{i}\in X$, sequences $(s_{k}^{(i,j)})_{k}$ and $(t_{k}^{(i)})_{k}$ in $\mathbb{N}$ have been constructed for $i=1,2,\dotsc,n$ and $j=1,\dotsc,i$ such that
  1. (1)
    $d(x_{i},y_{i})<\frac{1}{i}$ for $i=1,2,\dotsc,n$;
  2. (2)
    for $i=2,\dotsc,n$ and $j=1,\dotsc,i-1$, $(s_{k}^{(i,j)})_{k}$ is a subsequence of $(s_{k}^{(i-1,j)})_{k}$, and $(s_{k}^{(i,i)})_{k}$ is a subsequence of $(t_{k}^{(i-1)})_{k}$;
  3. (3)
    for $i=1,2,\dotsc,n$ and $j=1,\dotsc,i$,
    $\lim_{k\to\infty}d(T^{s_{k}^{(i,j)}}{x_{i}},\mathbf{0})=0\text{ and }\lim_{k\to\infty}p_{m}(T^{t_{k}^{(i)}}{x_{i}})=\infty;$
  4. (4)
    $\operatorname{span}\{x_{1},x_{2},\dotsc,x_{n}\}$ is an irregular manifold.
By the Claim,
$X_{n+1}:=\bigcap_{j=1}^{n}P(s_{k}^{(n,j)})\cap P(t_{k}^{(n)})\cap D$
is also a dense $G_{\delta}$ subset of $X$. Pick $x_{n+1}\in X_{n+1}$ with $d(x_{n+1},y_{n+1})<\frac{1}{n+1}$. For $j=1,\dotsc,n$, there exists a subsequence $(s_{k}^{(n+1,j)})_{k}$ of $(s_{k}^{(n,j)})_{k}$ such that
$\lim_{k\to\infty}d(T^{s_{k}^{(n+1,j)}}{x_{n+1}},\mathbf{0})=0,$
a subsequence $(s_{k}^{(n+1,n+1)})_{k}$ of $(t_{k}^{(n)})_{k}$ such that
$\lim_{k\to\infty}d(T^{s_{k}^{(n+1,n+1)}}{x_{n+1}},\mathbf{0})=0,$
and an increasing sequence $(t_{k}^{(n+1)})_{k}$ in $\mathbb{N}$ such that
$\lim_{k\to\infty}p_{m}(T^{t_{k}^{(n+1)}}{x_{n+1}})=\infty.$
For any $\sum_{\ell=1}^{n+1}\alpha_{\ell}x_{\ell}\in\operatorname{span}\{x_{1},\dotsc,x_{n},x_{n+1}\}\setminus\{\mathbf{0}\}$,
$\liminf_{k\to\infty}d\biggl{(}T^{s^{(n+1,1)}_{k}}\biggl{(}\sum_{\ell=1}^{n+1}\alpha_{\ell}x_{\ell}\biggl{)},\mathbf{0}\biggl{)}$ $\leq\sum_{\ell=1}^{n+1}(|\alpha_{\ell}|+1)\lim_{k\to\infty}d(T^{s^{(n+1,1)}_{k}}x_{\ell},\mathbf{0})=0.$
Let $i=\min\{\ell\in\{1,\dotsc,n+1\}\colon\alpha_{\ell}\neq 0\}$. If $i<n+1$, then
$\limsup_{k\to\infty}p_{m}\biggl{(}T^{s^{(n+1,i+1)}_{k}}\biggl{(}\sum_{\ell=i}^{n+1}\alpha_{\ell}x_{\ell}\biggl{)}\bigg{)}$ $\geq|\alpha_{i}|\lim_{k\to\infty}p_{m}(T^{s^{(n+1,i+1)}_{k}}x_{i})$
$\qquad-\sum_{\ell=i+1}^{n+1}|\alpha_{\ell}|\limsup_{k\to\infty}p_{m}(T^{s^{(n+1,i+1)}_{k}}x_{\ell})$
$\geq\infty-\sum_{\ell=i+1}^{n+1}|\alpha_{\ell}|2^{m}\limsup_{k\to\infty}d(T^{s^{(n+1,i+1)}_{k}}x_{\ell},\mathbf{0})$
$=\infty-0=\infty.$
If $i=n+1$, then by the construction one has
$\lim_{k\to\infty}p_{m}(T^{t_{k}^{(n+1)}}{x_{n+1}})=\infty.$
So $\operatorname{span}\{x_{1},\dotsc,x_{n},x_{n+1}\}$ is an irregular manifold. By induction, we obtain that the subspace $\operatorname{span}\{x_{i}\colon i\in\mathbb{N}\}$ is a dense irregular manifold. ∎
4. Mean Li-Yorke chaos
4.1. Mean equicontinuity, mean sensitivity and mean Li-Yorke chaos in topological dynamics
Let $(X,T)$ be a dynamical system with a metric $d$ on $X$. A pair $(x,y)\in X\times X$ is called mean asymptotic if
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0,$
mean proximal if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0,$
The mean asymptotic relation and mean proximal relation of $(X,T)$, denoted by $\operatorname{MAsym}(T)$ and $\operatorname{MProx}(T)$, are the set of all mean asymptotic pairs and mean proximal pairs respectively. For $x\in X$, the mean asymptotic cell and mean proximal cell of $x$ are defined by
$\operatorname{MAsym}(T,x)=\{y\in X\colon(x,y)\in\operatorname{MAsym}(T)\}$
and
$\operatorname{MProx}(T,x)=\{y\in X\colon(x,y)\in\operatorname{MProx}(T)\},$
respectively.
Similar to Lemma 3.1, we have the following $G_{\delta}$ sets.
Lemma 4.1.
Let $(X,T)$ be a dynamical system and $\delta>0$. Then
  1. (1)
    for every $x\in X$, the mean proximal cell of $x$ is a $G_{\delta}$ subset of $X$;
  2. (2)
    the mean proximal relation of $(X,T)$ is a $G_{\delta}$ subset of $X\times X$;
  3. (3)
    for every $x\in X$, the set
    $\biggl{\{}y\in X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta\biggr{\}}$
    is a $G_{\delta}$ subset of $X$;
  4. (4)
    the set
    $\biggl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta\biggr{\}}$
    is a $G_{\delta}$ subset of $X\times X$.
Proof.
(1) It follows from the following formula:
$\operatorname{MProx}(T,x)=\bigcap_{n=1}^{\infty}\biggl{\{}y\in X\colon\exists k>n\text{ s.t. }\frac{1}{k}\sum_{i=1}^{k}d(T^{i}x,T^{i}y)<\frac{1}{n}\biggr{\}}.$
(2), (3) and (4): These proofs are similar with (1). ∎
Following [LTY2015], we say that a dynamical system $(X,T)$ is mean equicontinuous if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x,y\in X$ with $d(x,y)<\delta$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)<\varepsilon.$
Lemma 4.2.
Let $(X,T)$ be a dynamical system. If $(X,T)$ is mean equicontinuous, then every proximal pair is mean asymptotic.
Proof.
Let $(x_{0},y_{0})\in X\times X$ be a proximal pair. As $(X,T)$ is mean equicontinuous, for every $\varepsilon>0$ there exists $\delta>0$ such that for any $x,y\in X$ with $d(x,y)<\delta$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)<\varepsilon.$
As $(x_{0},y_{0})$ is proximal, there exists $k\in\mathbb{N}$ such that $d(T^{k}x_{0},T^{k}y_{0})<\delta$. Then
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x_{0},T^{i}y_{0})=\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}(T^{k}x_{0}),T^{i}(T^{k}y_{0}))<\varepsilon.$
As $\varepsilon>0$ is arbitrary, $(x_{0},y_{0})$ is mean asymptotic. ∎
Following [LTY2015], we say that a dynamical system $(X,T)$ is mean sensitive if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ and
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta.$
The proof of the following result is similar to the one of Lemma 3.4, we leave the proof to the reader.
Lemma 4.3.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is mean sensitive;
  2. (2)
    there exists $\delta>0$ such that for every $x\in X$, the set
    $\biggl{\{}y\in X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta\biggr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    there exists $\delta>0$ such that the set
    $\biggl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta\biggr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
Let $(X,T)$ be a dynamical system. A pair $(x,y)\in X\times X$ is called mean Li-Yorke chaotic if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)>0,$
that is, $(x,y)$ is mean proximal but not mean asymptotic. A subset $K$ of $X$ is called mean Li-Yorke scrambled if any two distinct points $x,y\in K$ form a mean Li-Yorke chaotic pair. We say that the dynamical system $(X,T)$ is mean Li-Yorke chaotic if there exists an uncountable mean Li-Yorke scrambled subset of $X$.
For a given positive number $\delta$, a pair $(x,y)\in X\times X$ is called mean Li-Yorke $\delta$-chaotic if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta.$
Similarly, we can define mean Li-Yorke $\delta$-scrambled set and mean Li-Yorke $\delta$-chaos.
We say that a dynamical system $(X,T)$ is mean Li-Yorke sensitive if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that $(x,y)$ is mean Li-Yorke $\delta$-chaotic. It is clear that every mean Li-Yorke sensitive system is mean sensitive.
Similar to Proposition 3.5, according to Lemmas 4.1 and 4.3 we have the following result.
Proposition 4.4.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke $\delta$-scrambled;
  2. (2)
    the mean proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is mean sensitive.
From now on in this subsection, we fix a continuous pseudometric $\rho$ on $X$. We say that a dynamical system $(X,T)$ is mean $\rho$-extremely sensitive if for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ and
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,T^{i}y)=\infty.$
The proof of the following result is routine.
Lemma 4.5.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is mean $\rho$-extremely sensitive;
  2. (2)
    for every $x\in X$, the set
    $\biggl{\{}y\in X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,T^{i}y)=\infty\biggr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    the set
    $\biggl{\{}(x,y)\in X\times X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,T^{i}y)=\infty\biggr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
Let $(X,T)$ be a dynamical system. A pair $(x,y)\in X\times X$ is called mean Li-Yorke $\rho$-extremely chaotic if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,T^{i}y)=\infty.$
Similarly, we can define mean Li-Yorke $\rho$-extremely scrambled set, mean Li-Yorke $\rho$-extreme chaos and mean Li-Yorke $\rho$-extreme sensitivity.
Similar to Proposition 4.4, we have the following result.
Proposition 4.6.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke $\rho$-extremely scrambled;
  2. (2)
    the mean proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is mean $\rho$-extremely sensitive.
4.2. Mean Li-Yorke chaos for continuous endomorphisms of completely metrizable groups
In [BBP2020] the authors characterized of Li-Yorke chaos for linear operators on Banach spaces. In this subsection we consider the general setting of continuous endomorphisms of completely metrizable groups.
In this subsection, we fix a completely metrizable group $G$ and a left-invariant compatible metric $d$ on $G$.
Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called mean semi-irregular for $T$ if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)>0.$
The proof of the following result is similar to that of Lemma 3.9.
Lemma 4.7.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For every $x\in G$, $\operatorname{MAsym}(T,x)=x\cdot\operatorname{MAsym}(T,e)$, $\operatorname{MProx}(T,x)=x\cdot\operatorname{MProx}(T,e)$.
  2. (2)
    $\operatorname{MAsym}(T,e)$ is dense in $G$ if and only if $\operatorname{MAsym}(T)$ is dense in $G\times G$.
  3. (3)
    If $\operatorname{MAsym}(T,e)$ is residual in $G$, then $\operatorname{MAsym}(T)=G$.
  4. (4)
    $\operatorname{MProx}(T,e)$ is dense in $G$ if and only if $\operatorname{MProx}(T)$ is dense in $G\times G$.
  5. (5)
    For every $x,y\in G$, $(x,y)$ is a mean Li-Yorke chaotic pair if and only if $x^{-1}y$ is a mean semi-irregular point.
  6. (6)
    If $K$ is a mean Li-Yorke scrambled subset of $G$, then for any $x\in G$, $xK$ is also a mean Li-Yorke scrambled set. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of mean semi-irregular points.
  7. (7)
    For a given $\delta>0$, if $K$ is a mean Li-Yorke $\delta$-scrambled subset of $G$, then for any $x\in G$, $xK$ is also a mean Li-Yorke $\delta$-scrambled set.
Lemma 4.8.
Let $T\in\operatorname{End}(G)$. Then $\operatorname{MAsym}(T,e)$ is a $T$-invariant subgroup of $G$. If in addition the metric $d$ is bounded or the endomorphism $T$ is Lipschitz continuous, then for any increasing sequence $(n_{k})_{k}$ of positive integers,
$G(n_{k}):=\biggl{\{}x\in G\colon\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,e)=0\biggr{\}}$
is a $T$-invariant subgroup of $G$.
Proof.
For every $x,y\in G(n_{k})$, we have
$\limsup_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}(y^{-1}x),e)$ $=\limsup_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,T^{i}y)$
$\leq\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,e)+\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(e,T^{i}y)$
$=0.$
Then $G(n_{k})$ is a subgroup of $G$. Similarly, $\operatorname{MAsym}(T,e)$ is also a subgroup of $G$.
For every $x\in\operatorname{MAsym}(T,e)$,
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}(Tx),e)$ $=\lim_{n\to\infty}\frac{n+1}{n}\frac{1}{n+1}\biggl{(}\sum_{i=1}^{n+1}d(T^{i}x,e)-d(Tx,e)\biggr{)}$
$=\lim_{n\to\infty}\frac{1}{n+1}\sum_{i=1}^{n+1}d(T^{i}x,e)=0.$
Then $Tx\in\operatorname{MAsym}(T,e)$, which implies that $\operatorname{MAsym}(T,e)$ is $T$-invariant.
If in addition $d$ is bounded by $M$, then for every $x\in G(n_{k})$,
$\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}(Tx),e)$ $=\lim_{k\to\infty}\frac{1}{n_{k}}\biggl{(}\sum_{i=1}^{n_{k}}d(T^{i}x,e)+d(T^{n_{k}+1}x,e)-d(Tx,e)\biggr{)}$
$\leq\lim_{k\to\infty}\frac{1}{n_{k}}\biggl{(}\sum_{i=1}^{n_{k}}d(T^{i}x,e)+2M\biggr{)}=0.$
In this case, we have $Tx\in G(n_{k})$. Then $G(n_{k})$ is $T$-invariant.
If in addition $T$ is Lipschitz continuous, that is, there exists $L>0$ such that $d(Tx,Ty)\leq Ld(x,y)$ for any $x,y\in G$, then every $x\in G(n_{k})$,
$\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}(Tx),e)$ $=\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T(T^{i}x),Te)$
$\leq L\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,e)=0.$
In this case, we also have $Tx\in G(n_{k})$. Then $G(n_{k})$ is $T$-invariant. ∎
We have the following dichotomy of mean equicontinuity and mean sensitivity for continuous endomorphisms of completely metrizable groups.
Theorem 4.9.
Let $T\in\operatorname{End}(G)$. Then $T$ is either mean equicontinuous or mean sensitive.
Proof.
Let $e$ be the identity of $G$. We have the following two cases:
Case 1: for any $\varepsilon>0$, there exists $\delta>0$ such that for every $x\in G$ with $d(x,e)<\delta$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)<\varepsilon.$
Let $x_{1},x_{2}\in G$. If $d(x_{1},x_{2})<\delta$, by the left-invariance of $d$ we have $d(x_{1},x_{2})=d(x_{2}^{-1}x_{1},e)<\delta$. Then
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x_{1},T^{i}x_{2})=\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}(x_{2}^{-1}x_{1}),e)<\varepsilon.$
This implies that $T$ is mean equicontinuous.
Case 2: there exists $\delta>0$ such that for every $\varepsilon>0$ there exists $y_{\varepsilon}\in G$ with $d(y_{\varepsilon},e)<\varepsilon$ such that
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}y_{\varepsilon},e)\geq\delta.$
For every $x\in G$ and $\varepsilon>0$, we have $d(xy_{\varepsilon},x)<\varepsilon$ and
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}(xy_{\varepsilon}),T^{i}x)=\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}y_{\varepsilon},e)\geq\delta.$
Thus $T$ is mean sensitive. ∎
Combining Lemma 4.2 and Theorem 4.9, we have the following corollary.
Corollary 4.10.
Let $T\in\operatorname{End}(G)$. If there exists a mean Li-Yorke chaotic pair, then $T$ is mean sensitive.
Proposition 4.11.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is mean sensitive;
  2. (2)
    there exists $\delta>0$ such that for any $x\in G$ and $\varepsilon>0$ there exists $y\in G$ with $d(x,y)<\varepsilon$ and
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)\geq\delta.$
  3. (3)
    there exists a sequence $(y_{k})_{k}$ in $G$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=e$ and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(T^{i}y_{k},e)>0.$
Proof.
The implications (1$\Rightarrow$(2$\Rightarrow$(3) are clear.
(3$\Rightarrow$(1) As $Te=e$, without loss of generality, we can require the sequence $(N_{k})_{k}$ is increasing. Let
$\delta=\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}d(T^{i}y_{k},e).$
For each $n\in\mathbb{N}$, let
$X_{n}=\biggl{\{}x\in G\colon\exists k>n\text{ s.t. }\frac{1}{k}\sum_{i=1}^{k}d(T^{i}x,e)>\frac{\delta}{2}-\frac{1}{n}\biggr{\}}.$
Then each $X_{n}$ is an open subset of $G$. We show that it is also dense in $G$. Let $U$ be a nonempty open subset of $G$ and pick $x\in U$. If
$\limsup_{k\to\infty}\frac{1}{k}\sum_{i=1}^{k}d(T^{i}x,e)\geq\frac{\delta}{2},$
then $x\in X_{n}$. Now assume that
$\limsup_{k\to\infty}\frac{1}{k}\sum_{i=1}^{k}d(T^{i}x,e)<\frac{\delta}{2},$
Then there exists $N\in\mathbb{N}$ such that for any $k\geq N$,
$\frac{1}{k}\sum_{i=1}^{k}d(T^{i}x,e)\leq\frac{\delta}{2}.$
As $\lim_{k\to\infty}y_{k}x=x\in U$, there exists $k_{0}\in\mathbb{N}$ such that $y_{k_{0}}x\in U$ and $N_{k_{0}}>\max\{N,n\}$. Then
$\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}d(T^{i}(y_{k_{0}}x),e)$ $\geq\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}d(T^{i}y_{k_{0}},e)-\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}d(T^{i}(y_{k_{0}}x),T^{i}y_{k_{0}})$
$\geq\delta-\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}d(T^{i}x,e)$
$\geq\delta-\frac{\delta}{2}=\frac{\delta}{2}.$
This implies that $y_{k_{0}}x\in X_{n}\cap U$. Therefore, the set
$\biggl{\{}x\in G\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)\geq\frac{\delta}{2}\biggr{\}}=\bigcap_{n=1}^{\infty}X_{n}$
is residual in $G$. By the proof of Theorem 4.9, $T$ is mean sensitive. ∎
Combining Theorem 4.9 and Proposition 4.11, we have the following corollary.
Corollary 4.12.
Let $T\in\operatorname{End}(G)$. Then $T$ is mean equicontinuous if and only if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x,y\in G$ with $d(x,y)<\delta$,
$\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,T^{i}y)<\varepsilon.$
Now we have the following characterization about the existence of a dense set of mean semi-irregular points for continuous endomorphisms of completely metrizable groups. Since the proof is similar to the one of Theorem 3.13, we leave it to the reader.
Theorem 4.13.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of mean semi-irregular points;
  2. (2)
    $T$ admits a residual set of mean semi-irregular points;
  3. (3)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke $\delta$-scrambled;
  4. (4)
    the mean proximal relation of $T$ is dense in $G\times G$ and $T$ is mean sensitive;
  5. (5)
    the mean proximal cell of $e$ is dense in $G$ and there exists $x\in G$ such that
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)>0.$
According to Theorem 4.13 (5), we have the following consequences.
Corollary 4.14.
Let $T\in\operatorname{End}(G)$. If the mean proximal cell of $e$ is dense in $G$ and $T$ has a non-trivial periodic point, then it has a dense set of mean semi-irregular points.
Similar to the proof of Theorem 3.15, applying Theorem 4.13 we have the following characterization of mean Li-Yorke chaos for continuous endomorphisms of completely metrizable groups. Note that because we need Lemma 4.8, the metric $d$ on $G$ should be bounded or the $T$ should be Lipschitz continuous.
Theorem 4.15.
Let $T\in\operatorname{End}(G)$. If $d$ is a bounded left-invariant compatible metric on $G$ or $T$ is Lipschitz continuous, then the following assertions are equivalent:
  1. (1)
    $T$ admits a mean semi-irregular point;
  2. (2)
    $T$ admits a mean Li-Yorke chaotic pair;
  3. (3)
    $T$ is mean Li-Yorke chaotic;
  4. (4)
    $T$ is mean Li-Yorke sensitive;
  5. (5)
    the restriction of $T$ to some invariant closed $T$-invariant subgroup $\widetilde{G}$ has a dense set of mean semi-irregular points.
Similar to Proposition 3.17, we have the following result.
Proposition 4.16.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a residual mean Li-Yorke scrambled set;
  2. (2)
    every point in $G\setminus\{e\}$ is mean semi-irregular for $T$;
  3. (3)
    $X$ is a mean Li-Yorke scrambled set for $T$.
Example 4.17.
Let $\mathbb{S}(\mathbb{Z}),\sigma,T,A$ as in Example 3.19. As the mean proximal relation contains the asymptotic relation and $\sigma$ is a fixed point for $T$, according to Theorem 4.13 $(\mathbb{S}(\mathbb{Z}),T)$ is densely mean Li-Yorke chaotic for any left-invariant compatible metric on $\mathbb{S}(\mathbb{Z})$.
Example 4.18.
Let $H_{+}(\mathbb{R}),\sigma,T,A$ as in Example 3.20. Similar to the idea in Example 4.17, one has $(H_{+}(\mathbb{R}),T)$ is densely mean Li-Yorke chaotic for any left-invariant compatible metric on $H_{+}(\mathbb{R})$.
4.3. Mean Li-Yorke extreme chaos for continuous endomorphisms of completely metrizable groups
With the same as in the previous subsection, we fix a completely metrizable group $G$ and a left-invariant compatible metric $d$ on $G$. Similar to seminorms for Fréchet space, in this subsection we consider mean Li-Yorke chaos with respect to a left-invariant continuous pseudometric for metrizable groups.
In this subsection, we also fix a left-invariant continuous pseudometric $\rho$ on $G$. Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called mean irregular for $T$ if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,e)=0\qquad\text{and}\qquad\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,e)=\infty.$
The proofs of the following three results are similar to that of Lemma 3.9, Propositions 4.11 and 3.23 respectively.
Lemma 4.19.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For every $x,y\in G$, $(x,y)$ is a mean Li-Yorke extremely chaotic pair if and only if $x^{-1}y$ is a mean irregular point.
  2. (2)
    If $K$ is a mean Li-Yorke extremely scrambled subset of $G$, then for any $x\in G$, $xK$ is also a mean Li-Yorke extremely scrambled set. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of mean irregular points.
Proposition 4.20.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is mean extremely sensitive;
  2. (2)
    for any $x\in G$ and $\varepsilon>0$ there exists $y\in G$ with $d(x,y)<\varepsilon$ and
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\rho(T^{i}x,T^{i}y)=\infty.$
  3. (3)
    there exists a sequence $(y_{k})_{k}$ in $G$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=e$ and
    $\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\rho(T^{i}y_{k},e)\geq k,\ \forall k\in\mathbb{N}.$
Proposition 4.21.
Let $T\in\operatorname{End}(G)$. If there exists a mean irregular point, then $T$ is mean extremely sensitive.
Similar to Theorem 4.13, we have the following characterization about the existence of a dense set of mean irregular points for continuous endomorphisms of completely metrizable groups.
Theorem 4.22.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of mean irregular points;
  2. (2)
    $T$ admits a residual set of mean irregular points;
  3. (3)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke extremely scrambled;
  4. (4)
    the mean proximal relation of $T$ is dense in $G\times G$ and $T$ is mean extremely sensitive.
Applying Theorem 4.22 we have the following characterization of mean Li-Yorke extreme chaos for continuous endomorphisms of completely metrizable groups.
Theorem 4.23.
Let $T\in\operatorname{End}(G)$. If $d$ is a bounded left-invariant compatible metric on $G$ or $T$ is Lipschitz continuous, then the following assertions are equivalent:
  1. (1)
    $T$ admits a mean irregular point;
  2. (2)
    $T$ admits a mean Li-Yorke extremely chaotic pair;
  3. (3)
    $T$ is mean Li-Yorke extremely chaotic;
  4. (4)
    $T$ is mean Li-Yorke extremely sensitive;
  5. (5)
    the restriction of $T$ to some invariant closed $T$-invariant subgroup $\widetilde{G}$ has a dense set of mean irregular points.
4.4. Mean Li-Yorke chaos for continuous linear operators on Fréchet spaces
In this subsection unless otherwise specified, $X$ denotes an arbitrary infinite-dimensional Fréchet space. The separating increasing sequence $(p_{j})_{j}$ of seminorms on $X$ and the metric $d$ on $X$ are as in subsection 3.4. Though Li-Yorke chaos for linear operators on Banach spaces has been characterized in [BBP2020], in this subsection we generalize the results to the setting of linear operators on Fréchet spaces by apply results in the subsections 4.2 and 4.3 to continuous linear operators of $X$. We will also obtain some new results utilizing the linear structure of $X$.
As the metric $d$ on $X$ is bounded and $d(\lambda x,\mathbf{0})\leq(1+|\lambda|)d(x,\mathbf{0})$ for any $\lambda\in\mathbb{K}$ and $x\in X$, following ideas in Lemmas 3.27 and 4.8, we have the following result.
Lemma 4.24.
Let $T\in L(X)$ and $(n_{k})_{k}$ be an increasing sequence of positive integers. Then
$\biggl{\{}x\in X\colon\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,\mathbf{0})=0\biggr{\}}$
is a $T$-invariant subspace of $X$.
With the same idea as in subsection 3.4, we say that $T\in L(X)$ is mean $m$-extremely sensitive if it is mean extremely sensitive with respect to the seminorm $p_{m}$.
Proposition 4.25.
Let $T\in L(X)$ and $m\in\mathbb{N}$. Then the following assertions are equivalent:
  1. (1)
    $T$ is mean $m$-extremely sensitive;
  2. (2)
    there exists a vector $x\in X$ such that
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}p_{m}(T^{i}x)=\infty.$
  3. (3)
    there exists a bounded sequence $(y_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that
    $\sup_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}p_{m}(T^{i}y_{k})=\infty;$
  4. (4)
    there exists a sequence $(y_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=\mathbf{0}$ and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}p_{m}(T^{i}y_{k})>0.$
Proof.
The implication (1$\Rightarrow$(2$\Rightarrow$(3) are clear.
(3$\Rightarrow$(4) For each $k\in\mathbb{N}$, let $a_{k}=\frac{1}{N_{k}}\sum\limits_{i=1}^{N_{k}}p_{m}(T^{i}y_{k})$. Without loss of generality, we can assume that $(a_{k})_{k}$ is increasing. For each $k\in\mathbb{N}$, let $x_{k}=\frac{y_{k}}{a_{k}}$. Then $\lim_{k\to\infty}x_{k}=\mathbf{0}$ and
$\inf_{k}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}p_{m}(T^{i}x_{k})\geq 1.$
(4$\Rightarrow$(1) Let
$\delta=\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}p_{m}(T^{i}y_{k}).$
For each $n\in\mathbb{N}$, let
$X_{n}=\biggl{\{}x\in X\colon\exists k>n\text{ s.t. }\frac{1}{k}\sum_{i=1}^{k}p_{m}(T^{i}x)>n\biggr{\}}.$
It is clear that $X_{n}$ is open. We show that $X_{n}$ is dense. Let $U$ be a nonempty open subset of $X$ and pick $y\in X$. If
$\limsup_{k\to\infty}\frac{1}{k}\sum_{i=1}^{k}p_{m}(T^{i}y)=\infty,$
then $y\in X_{n}$. Now assume that
$\limsup_{k\to\infty}\frac{1}{k}\sum_{i=1}^{k}p_{m}(T^{i}y)=M<\infty.$
Then there exists $N\in\mathbb{N}$ such that for any $k\geq N$,
$\frac{1}{k}\sum_{i=1}^{k}p_{m}(T^{i}y)\leq 2M.$
There exists $k_{0}\in\mathbb{N}$ such that $y+\frac{2M+2n}{\delta}y_{k_{0}}\in U$ and $N_{k_{0}}>\max\{N,n\}$. Then
$\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}p_{m}(T^{i}(y+\tfrac{M+2n}{\delta}y_{k_{0}}))$ $\geq\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}p_{m}(T^{i}(\tfrac{2M+2n}{\delta}y_{k_{0}}))-\frac{1}{N_{k_{0}}}\sum_{i=1}^{N_{k_{0}}}p_{m}(T^{i}y)$
$\geq\tfrac{2M+2n}{\delta}\cdot\delta-2M>n$
This implies that $y+\frac{2M+2n}{\delta}x_{j_{0}}\in X_{n}$. Therefore, the set
$\biggl{\{}x\in X\colon\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}p_{m}(T^{i}x)=\infty\biggr{\}}=\bigcap_{n=1}^{\infty}X_{n}$
is residual in $X$, which implies that $T$ is mean $m$-extremely sensitive. ∎
Proposition 4.26.
Let $T\in L(X)$. If $T$ is mean sensitive then $T$ is mean $m$-extremely sensitive for some $m\in\mathbb{N}$.
Proof.
As $T$ is mean sensitive, there exists $\delta_{1}>0$ such that for every $k\in\mathbb{N}$ there exists $y_{k}\in X$ such that $d(y_{k},\mathbf{0})<\frac{1}{k}$ and
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}y_{k},\mathbf{0})>\delta_{1}.$
Note that $d$ is bounded by $1$, for every $x\in X$ and $t\in(0,1)$,
$\sum_{i=1}^{n}d(T^{i}x,\mathbf{0})\leq t(n-\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon d(T^{i}x,\mathbf{0})>t\}))+\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon d(T^{i}x,\mathbf{0})>t\}),$
then
$\frac{1}{n}\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon d(T^{i}x,\mathbf{0})>t\})\geq\frac{1}{1-t}\Bigl{(}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,\mathbf{0})-t\Bigr{)}.$
Pick $\delta_{2}>0$ such that for any $k\in\mathbb{N}$,
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{i}y_{k},\mathbf{0})>\delta_{2}\})\geq\delta_{2}.$
For every $x\in X$ and $m\in\mathbb{N}$,
$d(x,\mathbf{0})$ $=\sum_{j=1}^{\infty}\frac{1}{2^{j}}\min(1,p_{j}(x))\leq\sum_{j=1}^{m}\frac{1}{2^{j}}p_{j}(x)+\sum_{j={m+1}}^{\infty}\frac{1}{2^{j}}$
$\leq\sum_{j=1}^{m}\frac{1}{2^{j}}p_{m}(x)+\frac{1}{2^{m}}<p_{m}(x)+\frac{1}{2^{m}}.$
Pick $\delta>0$ and $m\in\mathbb{N}$ such that for any $x\in X$, $d(x,\mathbf{0})>\delta_{2}$ implies $p_{m}(x)>\frac{\delta}{\delta_{2}}$. Then for any $k\in\mathbb{N}$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}p_{m}(T^{i}y_{k})\geq\tfrac{\delta}{\delta_{2}}\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{i}y_{k},\mathbf{0})>\delta_{2}\})\geq\tfrac{\delta}{\delta_{2}}\delta_{2}=\delta.$
By Proposition 4.25, $T$ is mean $m$-extremely sensitive. ∎
In [BBP2020] the authors introduced the concepts of absolutely mean semi-irregular vector and absolutely mean irregular vector for linear operator of Banach space. It can be extend to Fréchet space with respect to the seminorm $p_{m}$ naturally. With the same idea as in subsection 3.4, we can define mean Li-Yorke $m$-extreme scrambled set, mean Li-Yorke $m$-extreme chaos and mean Li-Yorke $m$-extreme sensitivity.
Now we have the following characterization about the existence of a dense set of absolutely mean $m$-irregular vectors for continuous linear operators on Fréchet spaces. Since the proof is similar to the one of Theorem 3.32, we leave it to the reader.
Theorem 4.27.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of absolutely mean semi-irregular vectors;
  2. (2)
    $T$ admits a residual set of absolutely mean $m$-irregular vectors for some $m\in\mathbb{N}$;
  3. (3)
    there exists $m\in\mathbb{N}$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke $m$-extremely scrambled;
  4. (4)
    the mean proximal relation of $T$ is dense in $X\times X$ and $T$ is mean $m$-extremely sensitive for some $m\in\mathbb{N}$;
  5. (5)
    the mean proximal cell of $\mathbf{0}$ is dense in $X$ and there exists $x\in X$ such that
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}d(T^{i}x,\mathbf{0})>0.$
Similar to the proof of Theorem 3.33, applying Theorem 4.27 we get the following characterization of mean Li-Yorke $m$-extreme chaos for continuous linear operators on Fréchet spaces.
Theorem 4.28.
Let $T\in L(X)$ and $m\in\mathbb{N}$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits an absolutely mean $m$-irregular vector;
  2. (2)
    $T$ admits a mean Li-Yorke $m$-extremely chaotic pair;
  3. (3)
    $T$ is mean Li-Yorke $m$-extremely chaotic;
  4. (4)
    $T$ is mean Li-Yorke $m$-extremely sensitive;
  5. (5)
    the restriction of $T$ to some invariant closed $T$-invariant subspace $\widetilde{X}$ has a dense set of absolutely mean $m$-irregular vectors.
It is easy to see that if a vector is absolutely mean $m$-irregular for some $m\in\mathbb{N}$ then it is also absolutely mean semi-irregular. The following result reveals that any absolutely mean semi-irregular vector is an accumulation point of the set of absolutely mean $m$-irregular vectors for some $m\in\mathbb{N}$.
Proposition 4.29.
Let $T\in L(X)$. Then the set of absolutely mean $m$-irregular vectors for some $m\in\mathbb{N}$ is dense in the set of absolutely mean semi-irregular vectors.
Proof.
Let $y$ be an absolutely mean semi-irregular vector. There exists an increasing sequence $(n_{k})_{k}$ of positive integers such that
$\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}y,\mathbf{0})=0.$
By Lemma 4.24,
$X_{0}:=\biggl{\{}x\in X\colon\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}d(T^{i}x,\mathbf{0})=0\biggr{\}}$
is a $T$-invariant subspace of $X$. Let $\widetilde{X}$ be the closure of $X_{0}$. Then $\widetilde{X}$ is a $T$-invariant closed subspace of $X$. As the mean proximal cell of $e$ contains $X_{0}$, the mean proximal relation of $T|_{\widetilde{X}}$ is dense in $\widetilde{X}\times\widetilde{X}$. As $y\in\widetilde{X}$, $(y,e)$ is a mean Li-Yorke chaotic pair, then by Corollary 4.10, $T|_{\widetilde{X}}$ is mean sensitive. By Proposition 4.25, there exists $m\in\mathbb{N}$ and a vector $x\in\widetilde{X}$ with
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}p_{m}(T^{i}x)=\infty.$
Applying Theorem 4.27 to $T|_{\widetilde{X}}$ and $m$, the set of absolutely mean $m$-irregular vectors is residual in $\widetilde{X}$. Then $y$ is an accumulation point of the set of absolutely mean $m$-irregular vectors. ∎
Let $T\in L(X)$ and $m\in\mathbb{N}$. A vector subspace $Y$ of $X$ is called a absolutely mean $m$-irregular manifold for $T$ if every vector $y\in Y\setminus\{0\}$ is absolutely mean $m$-irregular for $T$. The following result was proved in [BBP2020]*Theorem 29 in the case of separable Banach space. Since the proof is similar to that of Theorem 3.36, we leave it to the reader.
Proposition 4.30.
Let $X$ be a separable Fréchet space and $T\in L(X)$. If the mean asymptotic cell of $\mathbf{0}$ is dense in $X$ then either the mean asymptotic cell of $\mathbf{0}$ is $X$ or $T$ admits a dense absolutely mean $m$-irregular manifold for some $m\in\mathbb{N}$.
4.5. Mean Li-Yorke chaos for continuous linear operators on Banach spaces
In this subsection, we consider continuous linear operators on Banach spaces. Let $(X,\lVert\cdot\rVert)$ be a Banach space. The usual metric induced by the norm is
$d_{B}(x,y)=\lVert x-y\rVert,\quad\forall x,y\in X.$
[…]
As every continuous linear operator of a Banach space is Lipschitz continuous, following ideas in Lemmas 4.8 and 4.24, we have the following result.
Lemma 4.31.
Let $(X,\lVert\cdot\rVert)$ be a Banach space, $T\in L(X)$ and $(n_{k})_{k}$ be an increasing sequence of positive integers. Then
$X_{0}:=\biggl{\{}x\in X\colon\lim_{k\to\infty}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\lVert T^{i}x\rVert=0\biggr{\}}$
is a $T$-invariant subspace of $X$.
Following [LH2015], we say that $T$ is absolutely Cesàro bounded if there exists a constant $C>0$ such that
$\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\|T^{i}x\|\leq C\|x\|$
for all $x\in X$. We have the following characterization of mean equicontinuity of continuous linear operators on Banach space, which can be regarded as a mean version of the Banach-Steinhaus Theorem.
Theorem 4.32.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is mean equicontinuous;
  2. (2)
    for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x\in X$ with $\lVert x\lVert<\delta$,
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert<\varepsilon;$
  3. (3)
    $T$ is absolutely Cesàro bounded;
  4. (4)
    for every $x\in X$,
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert<\infty.$
Proof.
(1$\Leftrightarrow$(2) It follows from Corollary 4.12.
(2$\Rightarrow$(3) There exists $\delta_{1}>0$ such that for every $x\in X$ with $\lVert x\lVert<\delta_{1}$,
$\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert<1.$
Let $C=\frac{2}{\delta_{1}}$. For every $x\in X$, $\bigl{\lVert}\frac{\delta_{1}x}{2\lVert x\rVert}\bigr{\rVert}<\delta_{1}$, and then
$\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert=\frac{2\lVert x\rVert}{\delta_{1}}\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\Bigl{\lVert}T^{i}\Bigl{(}\frac{\delta_{1}x}{2\lVert x\rVert}\Bigr{)}\Bigr{\rVert}<C\lVert x\rVert.$
This shows that $T$ is absolutely Cesàro bounded.
The implication (3$\Rightarrow$(4) is clear.
(4$\Rightarrow$(2) For each $N\in\mathbb{N}$, let
$X_{N}=\biggl{\{}x\in X\colon\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert\leq N\biggr{\}}.$
It is clear that $X_{n}$ is closed and $X=\bigcup_{n=1}^{\infty}X_{n}$. By the Baire category theorem, there exists $N_{0}\in\mathbb{N}$ such that the interior of $X_{N_{0}}$ is not empty. Then $X_{N_{0}}-X_{N_{0}}$ is a neighborhood of $\mathbf{0}$, that is, there exists $\delta_{0}>0$ such that $\{x\in X\colon\lVert x\rVert<\delta_{0}\}\subset X_{N_{0}}-X_{N_{0}}$. For any $x\in X$ with $\lVert x\rVert<\delta_{0}$, there exists $x_{1},x_{2}\in X_{N_{0}}$ such that $x=x_{1}-x_{2}$ and then
$\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\lVert$ $=\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}(x_{1}-x_{2})\lVert$
$\leq\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}(x_{1})\lVert+\sup_{n\in\mathbb{N}}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}(x_{2})\lVert$
$\leq N_{0}+N_{0}=2N_{0}.$
Thus (2) follows from the linearity of $T$. ∎
Combining Propositions 4.11 and 4.25, and Theorem 4.32, we have the following equivalent conditions for mean sensitive operators on Banach spaces.
Proposition 4.33.
Let $X$ be a Banach space and $T\in L(X)$.
  1. (1)
    $T$ is mean sensitive;
  2. (2)
    $T$ is not absolutely Cesàro bounded;
  3. (3)
    $T$ is mean extremely sensitive.
Now we have the following characterization about the existence of a dense set of mean irregular vectors for continuous linear operators on Banach spaces, which slightly strengthens [BBP2020]*Theorems 17 and 22. Since the proof is similar to the one of Theorem 3.32, we leave it to the reader.
Theorem 4.34.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of absolutely mean semi-irregular vectors;
  2. (2)
    $T$ admits a residual set of absolutely mean irregular vectors;
  3. (3)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is mean Li-Yorke extremely scrambled;
  4. (4)
    the mean proximal relation of $T$ is dense in $X\times X$ and $T$ is mean sensitive;
  5. (5)
    the mean proximal cell of $\mathbf{0}$ is dense in $X$ and there exists $x\in X$ such that
    $\limsup_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\rVert>0.$
According to Theorem 4.34 (5), we have the following consequences.
Corollary 4.35.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Assume that the mean proximal cell of $\mathbf{0}$ is dense in $X$. If $T$ has a non-trivial periodic point or an eigenvalue $\lambda$ with $|\lambda|\geq 1$, then it has a residual set of absolutely mean irregular vectors.
The following result characterizes mean Li-Yorke chaos for continuous linear operators on Banach spaces, which is essentially contained in [BBP2020]*Theorem 5. Here we can apply Thereon 4.34 to get a direct proof.
Theorem 4.36.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits an absolutely mean semi-irregular vector;
  2. (2)
    $T$ admits a mean Li-Yorke chaotic pair;
  3. (3)
    $T$ admits an absolutely mean irregular vector;
  4. (4)
    $T$ is mean Li-Yorke extremely chaotic;
  5. (5)
    $T$ is mean Li-Yorke extremely sensitive;
  6. (6)
    the restriction of $T$ to some invariant closed $T$-invariant subspace $\widetilde{X}$ has a residual set of absolutely mean irregular vectors.
Corollary 4.37 ([BBP2020]*Theorem 7).
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the set of absolutely mean irregular vectors is dense in the set of absolutely mean semi-irregular vectors.
Following [BBP2020], we say that $T\in L(X)$ satisfies the mean Li-Yorke chaos criterion if there exists a subset $X_{0}$ of $X$ with the following properties:
  1. (1)
    for every $x\in X_{0}$, $\liminf\limits_{n\to\infty}\frac{1}{n}\sum\limits_{i=1}^{n}\lVert T^{i}x\rVert=0$;
  2. (2)
    there exists a sequence $(y_{k})_{k}$ in $\overline{\operatorname{span}(X_{0})}$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that for every $k\in\mathbb{N}$,
    $\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\lVert T^{i}y_{k}\rVert\geq k\lVert y_{k}\rVert.$
Using the ideas developed in this subsection, we can give a new proof to the following result. Since the proof is similar to the one of Proposition 3.35, we leave it to the reader.
Proposition 4.38 ([BBP2020]*Theorem 9).
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then $T$ is mean Li-Yorke chaotic if and only if it satisfies the mean Li-Yorke chaos criterion.
Remark 4.39.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. As a Fréchet space, $X$ has the metric
$d_{F}(x,y)=\min(\lVert x-y\rVert,1),\quad\forall x,y\in X.$
It should be noticed that mean Li-Yorke chaos with respect to the metric $d_{B}$ may be different from the one with respect to the metric $d_{F}$.
In [BBPW2018]*Theorem 25, the authors constructed a continuous linear operator $T$ of $X=c_{0}(\mathbb{N})$ or $X=\ell^{p}(\mathbb{N})$ for some $1\leq p<\infty$ such that $T$ is distributionally chaotic of type $1$ (see the next section for the definition) and for all $x\in X\setminus\{0\}$,
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\lVert T^{i}x\rVert=\infty.$
Then it is easy to verify that $T$ is mean Li-Yorke chaotic with respect to $d_{F}$, but not with respect to $d_{B}$.
5. Distributional chaos
5.1. Mean-L-stability, mean-L-unstability and distributional chaos in topological dynamics
Let $(X,T)$ be a dynamical system with a metric $d$ on $X$. A pair $(x,y)\in X\times X$ is called distributionally asymptotic if for every $\varepsilon>0$,
$\operatorname{\mathrm{dens}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)<\varepsilon\})=1,$
distributionally proximal if for every $\varepsilon>0$,
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)<\varepsilon\})=1.$
The distributionally asymptotic relation and distributionally proximal relation of $(X,T)$, denoted by $\operatorname{DAsym}(T)$ and $\operatorname{DProx}(T)$, are the set of all distributionally asymptotic pairs and distributionally proximal pairs respectively. For $x\in X$, the distributionally asymptotic cell and distributionally proximal cell of $x$ are defined by
$\operatorname{DAsym}(T,x)=\{y\in X\colon(x,y)\in\operatorname{DAsym}(T)\}$
and
$\operatorname{DProx}(T,x)=\{y\in X\colon(x,y)\in\operatorname{DProx}(T)\},$
respectively.
Similar to Lemma 3.1, we have the following $G_{\delta}$ sets.
Lemma 5.1.
Let $(X,T)$ be a dynamical system and $\delta,\eta>0$. Then
  1. (1)
    for every $x\in X$, the distributionally proximal cell of $x$ is a $G_{\delta}$ subset of $X$;
  2. (2)
    the distributionally proximal relation is a $G_{\delta}$ subset of $X\times X$;
  3. (3)
    for every $x\in X$, the set
    $\bigl{\{}y\in X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\eta\bigr{\}}$
    is a $G_{\delta}$ subset of $X$;
  4. (4)
    the set
    $\bigl{\{}(x,y)\in X\times X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\eta\bigr{\}}$
    is a $G_{\delta}$ subset of $X\times X$.
Proof.
(1) follows from the following formula:
[…]
$\operatorname{\mathrm{card}}\bigl{(}\{1\leq i\leq k\colon d(T^{i}x,T^{i}y)<\tfrac{1}{n}\}\bigr{)}>k(1-\tfrac{1}{n})\bigr{\}}.$
(2), (3) and (4): These proofs are similar with (1). ∎
Following [F1951], we say that a dynamical system $(X,T)$ is stable in the mean in the sense of Lyapunov (or mean-L-stable briefly) if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x,y\in X$ with $d(x,y)<\delta$,
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)\geq\varepsilon\})<\varepsilon.$
Lemma 5.2.
Let $(X,T)$ be a dynamical system. If $(X,T)$ is mean-L-stable, then every proximal pair $(x,y)\in X\times X$ is distributionally asymptotic.
Proof.
Let $(x_{0},y_{0})\in X\times X$ be a proximal pair. As $(X,T)$ is mean-L-stable, for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x,y\in X$ with $d(x,y)<\delta$,
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)\geq\varepsilon\})<\varepsilon.$
As $(x_{0},y_{0})$ is proximal, there exists $k\in\mathbb{N}$ such that $d(T^{k}x_{0},T^{k}y_{0})<\delta$. Then
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x_{0},T^{n}y_{0})=\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}(T^{k}x_{0}),T^{n}(T^{k}y_{0}))\geq\varepsilon\})<\varepsilon.$
As $\varepsilon>0$ is arbitrary, $(x_{0},y_{0})$ is distributionally asymptotic. ∎
We say that a dynamical system $(X,T)$ is unstable in the mean in the sense of Lyapunov (or mean-L-unstable briefly) if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\delta.$
The proof of the following result is similar to the one of Lemma 3.4. We leave the proof to the reader.
Lemma 5.3.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is mean-L-unstable;
  2. (2)
    there exists $\delta>0$ such that for every $x\in X$, the set
    $\bigl{\{}y\in X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\delta\bigr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    there exists $\delta>0$ such that the set
    $\bigl{\{}(x,y)\in X\times X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\delta\bigr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
Let $(X,T)$ be a dynamical system. For a given $\delta>0$, a pair $(x,y)\in X\times X$ is called distributionally $\delta$-chaotic of type $2$ if $(x,y)$ is distributionally proximal and
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})\geq\delta.$
A subset $K$ of $X$ is called distributionally scrambled of type $2$ if any two distinct points $x,y\in K$ form a distributionally $\delta$-chaotic pair of type $2$ for some $\delta=\delta(x,y)>0$. We say that the dynamical system $(X,T)$ is distributionally chaotic of type $2$ if there exists an uncountable distributionally scrambled subset of type $2$.
Similarly, we can define distributionally $\delta$-scrambled set of type $2$ and distributional $\delta$-chaos of type $2$.
The proof of the following result is analogous to that of Proposition 3.5.
Proposition 5.4.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $\delta$-scrambled of type $2$;
  2. (2)
    the distributionally proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is mean-L-unstable.
Remark 5.5.
If the metric $d$ is bounded, it is observed in [D2014] that a pair is mean Li-Yorke chaotic if and only if it is distributionally chaotic of type $2$, and in [LTY2015] that $(X,T)$ is mean-L-unstable if and only if it is mean sensitive.
For a given $\delta>0$, a pair $(x,y)\in X\times X$ is called distributionally $\delta$-chaotic of type $1$ if $(x,y)$ is distributionally proximal and
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})=1.$
A subset $K$ of $X$ is called distributionally scrambled of type $1$ if any two distinct points $x,y\in K$ form a distributionally $\delta$-chaotic pair of type $1$ for some $\delta=\delta(x,y)>0$. We say that the dynamical system $(X,T)$ is distributionally chaotic of type $1$ if there exists an uncountable distributionally scrambled subset of type $1$.
Similarly, we can define distributionally $\delta$-scrambled set of type $1$ and distributional $\delta$-chaos of type $1$.
The proof of the following result is analogous to that of Proposition 3.5.
Proposition 5.6.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $\delta$-scrambled of type $1$;
  2. (2)
    the distributionally proximal relation of $(X,T)$ is dense in $X\times X$, and there exists $\delta>0$ such that the set
    $\bigl{\{}(x,y)\in X\times X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})=1\bigr{\}}$
    is dense in $X\times X$.
From now on in this subsection, we fix a continuous pseudometric $\rho$ on $X$. We say that a dynamical system $(X,T)$ is $\rho$-extremely mean-L-unstable if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that for any $M>0$
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})\geq\delta.$
The proof of the following result is similar to the one of Lemma 3.4. We leave the proof to the reader.
Lemma 5.7.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    $(X,T)$ is $\rho$-extremely mean-L-unstable;
  2. (2)
    there exists $\delta>0$ such that for every $x\in X$, the set
    $\bigl{\{}y\in X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})\geq\delta,\ \forall M>0\bigr{\}}$
    is a dense $G_{\delta}$ subset of $X$;
  3. (3)
    there exists $\delta>0$ such that the set
    $\bigl{\{}(x,y)\in X\times X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})\geq\delta,\forall M>0\bigr{\}}$
    is a dense $G_{\delta}$ subset of $X\times X$.
For a given $\delta>0$, a pair $(x,y)\in X\times X$ is called distributionally $\rho$-extremely $\delta$-chaotic of type $2$ if $(x,y)$ is distributionally proximal and for any $M>0$
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})\geq\delta,$
and distributionally $\rho$-extremely chaotic of type $1$ if $(x,y)$ is distributionally proximal and for any $M>0$
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})=1.$
Similarly, we can define distributionally $\rho$-extremely $\delta$-scrambled set of type $2$ and distributional $\rho$-extreme $\delta$-chaos of type $2$, distributionally $\rho$-extremely scrambled set of type $1$ and distributional $\rho$-extreme chaos of type $1$.
The proof of the following two results are analogous to that of Proposition 3.5.
Proposition 5.8.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $\rho$-extremely $\delta$-scrambled of type $2$;
  2. (2)
    the distributionally proximal relation of $(X,T)$ is dense in $X\times X$ and $(X,T)$ is $\rho$-extremely mean-L-unstable.
Proposition 5.9.
Let $(X,T)$ be a dynamical system with $X$ being completely metrizable. Then the following assertions are equivalent:
  1. (1)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $\rho$-extremely scrambled of type $1$;
  2. (2)
    the distributionally proximal relation of $(X,T)$ is dense in $X\times X$, and the set
    $\bigl{\{}(x,y)\in X\times X\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\rho(T^{n}x,T^{n}y)>M\})=1,\ \forall M>0\bigr{\}}$
    is dense in $X\times X$.
Remark 5.10.
According to Lemma 2.1, we can give equivalent definitions of distributional chaos of type $1$ and $2$ by limits along subsets of $\mathbb{N}$.
5.2. Distributional chaos for continuous endomorphisms of completely metrizable groups
In [BBMP2011], [BBMP2013] and [BBPW2018] the authors characterized of distribitional Li-Yorke chaos for linear operators on Banach and Fréchet spaces. In this section, we will generalize some of the results to the setting of continuous endomorphisms of completely metrizable groups.
In this subsection, we fix a completely metrizable group $G$ and a left-invariant compatible metric $d$ on $G$. Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called distributional semi-irregular of type $2$ for $T$ if there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x=e$ and there exists a subset $B$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(B)>0$ such that $\liminf\limits_{B\ni n\to\infty}d(T^{n}x,e)>0$, and distributional semi-irregular of type $1$ for $T$ if there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x=e$ and there exists a subset $B$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(B)=1$ such that $\liminf\limits_{B\ni n\to\infty}d(T^{n}x,e)>0$.
The proof of the following result is similar to that of Lemma 3.9.
Lemma 5.11.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For every $x\in G$, $\operatorname{DAsym}(T,x)=x\cdot\operatorname{DAsym}(T,e)$, $\operatorname{DProx}(T,x)=x\cdot\operatorname{DProx}(T,e)$;
  2. (2)
    $\operatorname{DAsym}(T,e)$ is dense in $G$ if and only if $\operatorname{DAsym}(T)$ is dense in $G\times G$;
  3. (3)
    If $\operatorname{DAsym}(T,e)$ is residual in $G$, then $\operatorname{DAsym}(T,e)=G$;
  4. (4)
    $\operatorname{DProx}(T,e)$ is dense in $G$ if and only if $\operatorname{DProx}(T)$ is dense in $G\times G$;
  5. (5)
    For $i=1,2$ and every $x,y\in G$, $(x,y)$ is distributionally chaotic of type $i$ if and only if $x^{-1}y$ is a distributional semi-irregular point of type $i$.
  6. (6)
    For $i=1,2$, if a subset $K$ of $G$ is distributionally scrambled of type $i$, then for any $x\in G$, $xK$ is also distributionally scrambled of type $i$. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of distributional semi-irregular points of type $i$.
  7. (7)
    For $i=1,2$ and a given $\delta>0$, if a subset $K$ of $G$ is distributional $\delta$-scrambled type $i$, then for any $x\in G$, $xK$ is also distributional $\delta$-scrambled type $i$
Remark 5.12.
Let $T\in\operatorname{End}(G)$. If $(x,e)$ is distributional proximal, then there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x=e$. Let
$G_{0}=\Bigl{\{}y\in G\colon\lim\limits_{A\ni n\to\infty}T^{n}y=e\Bigr{\}}.$
It is clear that $G_{0}\subset\operatorname{DProx}(T,e)$. By Lemma 3.10, $G_{0}$ is a subgroup of $G$.
The proof of the following result is analogous that of Theorem 4.9.
Theorem 5.13.
Let $T\in\operatorname{End}(G)$. Then $T$ is either mean-L-stable or mean-L-unstable.
Combining Lemma 5.2 and Theorem  5.13, we have the following corollary.
Corollary 5.14.
Let $T\in\operatorname{End}(G)$. If there exists a proximal pair $(x,y)\in G$ and $\delta>0$ such that $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta)>0$, then $T$ is mean-L-unstable. In particular, if there exists a distributionally chaotic pair of type $2$, then $T$ is mean-L-unstable.
The proof of the following result is analogous that of Theorem 4.11.
Proposition 5.15.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is mean-L-unstable;
  2. (2)
    there exists $\delta>0$ such that for any $x\in G$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ and
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon d(T^{i}x,T^{i}y)>\delta\})\geq\delta.$
  3. (3)
    there exists $\delta>0$, a sequence $(y_{k})_{k}$ in $G$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=e$ and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon d(T^{i}y_{k},e)>\delta\})\geq\delta.$
Combining Theorem 5.13 and Proposition 5.15, we have the following corollary.
Corollary 5.16.
Let $T\in\operatorname{End}(G)$. Then $T$ is mean-L-stable if and only if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $x,y\in G$ with $d(x,y)<\delta$,
$\sup_{n\in\mathbb{N}}\frac{1}{n}\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon d(T^{i}x,T^{i}y)\geq\varepsilon\})<\varepsilon.$
Now we have the following characterization about the existence of a dense set of distributionally semi-irregular points for continuous endomorphisms of completely metrizable groups. Since the proof is similar to the one of Theorem 3.13, we leave it to the reader.
Theorem 5.17.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributional semi-irregular points of type $2$;
  2. (2)
    $T$ admits a residual set of distributional semi-irregular points of type $2$;
  3. (3)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributional $\delta$-scrambled of type $2$;
  4. (4)
    the distributionally proximal relation of $T$ is dense in $G\times G$ and $T$ is mean-L-unstable;
  5. (5)
    the distributionally proximal cell of $e$ is dense in $G$ and there exists $x\in G$ and $\delta>0$ such that
    $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,e)>\delta)>0.$
According to Theorem 5.17 (5), we have the following consequences.
Corollary 5.18.
Let $T\in\operatorname{End}(G)$. If the distributionally proximal cell of $e$ is dense in $G$ and $T$ has a non-trivial periodic point, then it has a dense set of distributionally semi-irregular points of type $2$.
Applying Theorem 5.17, we have the following characterization of distributional chaos of type $2$ for continuous endomorphisms of completely metrizable groups.
Theorem 5.19.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a distributional semi-irregular point of type $2$;
  2. (2)
    $T$ admits a distributionally chaotic pair of type $2$;
  3. (3)
    $T$ is distributionally chaotic of type $2$;
  4. (4)
    the restriction of $T$ to some invariant closed $T$-invariant subgroup $\widetilde{G}$ has a dense set of distributional semi-irregular points of type $2$.
Similar to Proposition 3.17, we have the following result.
Proposition 5.20.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a residual distributionally scrambled set of type $2$;
  2. (2)
    every point in $G\setminus\{e\}$ is distributional semi-irregular of type $2$ for $T$;
  3. (3)
    $X$ is a distributionally scrambled set of type $2$ for $T$.
Example 5.21.
Let $\mathbb{S}(\mathbb{Z}),\sigma,T,A$ as in Example 3.19. As the distributional proximal relation contains the asymptotic relation and $\sigma$ is a fixed point for $T$, according to Theorem 5.17 $(\mathbb{S}(\mathbb{Z}),T)$ is densely distributionally chaotic of type $2$ for any left-invariant compatible metric on $\mathbb{S}(\mathbb{Z})$.
Let $d$ be the precise left-invariant compatible metric as in Example 3.19. It is easy to show that there exists $\delta>0$ such that the set
$\bigl{\{}(x,y)\in\mathbb{S}(\mathbb{Z})\times\mathbb{S}(\mathbb{Z})\colon\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})=1\bigr{\}}$
is dense in $\mathbb{S}(\mathbb{Z})\times\mathbb{S}(\mathbb{Z})$. Then by Proposition 5.6 $(\mathbb{S}(\mathbb{Z}),T)$ is densely distributionally chaotic of type $1$ with respect to this metric $d$.
Example 5.22.
Let $H_{+}(\mathbb{R}),\sigma,T,A$ as in Example 3.20. Similar to the idea in Example 4.17, one has $(H_{+}(\mathbb{R}),T)$ is densely distributionally chaotic of type $2$ for any left-invariant compatible metric on $H_{+}(\mathbb{R})$, and densely distributionally chaotic of type $1$ with respect to the precise metric compatible $d$ as in Example 3.20.
5.3. Distributional extreme chaos for continuous endomorphisms of completely metrizable groups
With the same as in the previous subsection, we fix a completely metrizable group $G$ and a left-invariant compatible metric $d$ on $G$. Similar to seminorms in Fréchet spaces, in this subsection we fix a left-invariant continuous pseudometric $\rho$ on $G$.
Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called distributionally irregular of type $2$ if there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x=e$, and there exists a subset $B$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(B)>0$ such that $\lim\limits_{B\ni n\to\infty}\rho(T^{n}x,e)=\infty$, and distributionally irregular of type $1$ for $T$ if there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x=e$, and there exists a subset $B$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(B)=1$ such that such that $\lim\limits_{B\ni n\to\infty}\rho(T^{n}x,e)=\infty$.
The proofs of the following three results are similar to that of Lemma 3.9, Propositions 4.11 and 3.23 respectively.
Lemma 5.23.
Let $T\in\operatorname{End}(G)$.
  1. (1)
    For $i=1,2$, and every $x,y\in G$, $(x,y)$ is distributionally extremely chaotic of type $i$ if and only if $x^{-1}y$ is a distributionally irregular point of type $i$.
  2. (2)
    For $i=1,2$, if $K\subset G$ is a distributionally extremely scrambled subset of type $i$, then for any $x\in G$, $xK$ is also a distributionally extremely scrambled set of type $i$. In particular, for any $x\in K$, $x^{-1}K\setminus\{e\}$ consists of distributionally irregular points of type $i$.
Proposition 5.24.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is extremely mean-L-unstable;
  2. (2)
    there exists $\delta>0$ such that for any $x\in G$ and $\varepsilon,M>0$ there exists $y\in G$ with $d(x,y)<\varepsilon$ and
    $\sup_{n\in\mathbb{N}}\frac{1}{n}\operatorname{\mathrm{card}}(\{1\leq i\leq n\colon\rho(T^{i}x,T^{i}y)>M\})\geq\delta.$
  3. (3)
    there exists a sequence $(y_{k})_{k}$ in $G$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=e$ and
    $\inf_{k\in\mathbb{N}}\frac{1}{N_{k}}\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon\rho(T^{i}y_{k},e)>k\})>0.$
Proposition 5.25.
Let $T\in\operatorname{End}(G)$. If there exists a distributionally irregular point of type $2$, then $T$ is extremely mean-L-unstable.
Now we have the following characterization about the existence of a dense set of distributionally irregular points of type $2$ for continuous endomorphisms of completely metrizable groups. Since the proof is similar to the one of Theorem 3.24, we leave it to the reader.
Theorem 5.26.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributionally irregular points of type $2$;
  2. (2)
    $T$ admits a residual set of distributionally irregular points of type $2$;
  3. (3)
    there exists $\delta>0$ for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally extremely $\delta$-scrambled of type $2$;
  4. (4)
    the distributionally proximal relation of $T$ is dense in $G\times G$ and $T$ is extremely mean-L-unstable.
Applying Theorem 5.26, we have the following characterization of distributional extreme chaos of type $2$ for continuous endomorphisms of completely metrizable groups.
Theorem 5.27.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a distributionally irregular point of type $2$;
  2. (2)
    $T$ admits a distributionally extremely chaotic pair of type $2$;
  3. (3)
    $T$ is distributionally extremely chaotic of type $2$;
  4. (4)
    the restriction of $T$ to some invariant closed $T$-invariant subgroup $\widetilde{G}$ has a dense set of distributionally irregular points of type $2$.
Let $T\in\operatorname{End}(G)$. A point $x\in G$ is called distributionally unbounded if there exists a subset $B$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(B)=1$ such that $\lim\limits_{B\ni n\to\infty}\rho(T^{n}x,e)=\infty$. Then a point $x\in G$ is distributionally irregular of type $1$ if and only if if $(x,e)$ is distributionally proximal and $x$ is distributionally unbounded. For each $n\in X$, let
$G_{n}=\bigl{\{}x\in G\colon\exists k>n\text{ s.t. }\operatorname{\mathrm{card}}(\{1\leq i\leq k\colon\rho(T^{i}x,e)>n\})>k(1-\tfrac{1}{n})\bigr{\}}.$
It is clear that $G_{n}$ is open and $\bigcap_{n=1}^{\infty}G_{n}$ is the set of distributionally unbounded points.
Similar to Theorem 5.26, we have the following characterization about the existence of a dense set of distributionally irregular points of type $1$ for continuous endomorphisms of completely metrizable groups.
Theorem 5.28.
Let $T\in\operatorname{End}(G)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributionally irregular points of type $1$;
  2. (2)
    $T$ admits a residual set of distributionally irregular points of type $1$;
  3. (3)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $G$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally extremely scrambled of type $1$;
  4. (4)
    the distributionally proximal cell of $e$ is dense in $G$ and $T$ admits a dense set of distributionally unbounded points.
5.4. Distributional chaos for continuous linear operators on Fréchet spaces
Distributional chaos of type 1 for linear operators on Fréchet spaces has been characterized in [BBMP2013], and four types of distributional chaos of type for linear operators on Banach spaces are investigated in [BBPW2018]. It should be notcied that in [YLW21] Yin et al. studied distributional chaos for operators on Fréchet spaces with respect to the bounded metric. In this subsection using results in the subsections 5.2 and 5.3 we will consider distributional chaos of type 1 and 2 linear operators on Fréchet spaces which reprove some results in [BBMP2013] and generalize some results in [BBPW2018] to the setting of linear operators on Fréchet spaces.
In this subsection unless otherwise specified, $X$ denotes an arbitrary infinite-dimensional Fréchet space. The separating increasing sequence $(p_{j})_{j}$ of seminorms on $X$ and the metric $d$ on $X$ are as in subsection 3.4.
Let $T\in L(X)$, $m\in\mathbb{N}$ and $i=1,2$. A vector $x\in X$ is called distributionally $m$-irregular of type $i$ if it is distributionally irregular of type $i$ with respect to the seminorm $p_{m}$. A pair $(x,y)\in X\times X$ is called distributionally $m$-extremely $\delta$-chaotic if it is distributionally extremely $\delta$-scrambled with respect to the seminorm $p_{m}$. Similarly, we can define distributional $m$-extreme sensitivity, distributionally $m$-extremely $\delta$-scrambled set of type $i$ and distributional $m$-extreme chaos of type $i$.
Proposition 5.29.
Let $T\in L(X)$. Then $T$ is mean-L-unstable if and only if it is $m$-extremely mean-L-unstable for some $m\in\mathbb{N}$.
Proof.
If $T$ is $m$-extremely mean-L-unstable for some $m\in\mathbb{N}$, as $d(x,y)\geq\frac{1}{2^{m}}p_{m}(x,y)$ for all $x,y\in X$, then $T$ is mean-L-unstable. Now assume that $T$ is mean-L-unstable. By Proposition 5.15, there exists $\delta>0$, a sequence $(y_{k})_{k}$ in $X$ and a sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=\mathbf{0}$ and
$\sup_{k\in\mathbb{N}}\frac{1}{N_{k}}\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon d(T^{i}y_{k},\mathbf{0})>\delta\})\geq\delta.$
As $d(x,\mathbf{0})<p_{m}(x)+\frac{1}{2^{m}}$ for all $x\in X$ and $m\in\mathbb{N}$, there exists $\delta^{\prime}>0$ and $m\in\mathbb{N}$ such that
$\sup_{k\in\mathbb{N}}\frac{1}{N_{k}}\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon p_{m}(T^{i}y_{k})>\delta^{\prime}\})\geq\delta.$
For each $k\in\mathbb{N}$, let $x_{k}=\frac{\delta^{\prime}}{k}y_{k}$. Then $\lim_{k\to\infty}x_{k}=\mathbf{0}$ and
$\sup_{k\in\mathbb{N}}\frac{1}{N_{k}}\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon p_{m}(T^{i}x_{k})>k\})\geq\delta.$
Then by Proposition 5.24, $T$ is $m$-extremely mean-L-unstable. ∎
Combining Theorems 5.17 and 5.26 and Proposition 5.29, we have the following characterization of the existence of a dense set of distributionally semi-irregular points of type $2$ for continuous linear operators on Fréchet spaces.
Theorem 5.30.
Let $T\in L(X)$ and $m\in\mathbb{N}$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributionally semi-irregular points of type $2$;
  2. (2)
    $T$ admits a residual set of distributionally $m$-irregular points of type $2$ for some $m\in\mathbb{N}$;
  3. (3)
    there exists $\delta>0$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $m$-extremely $\delta$-scrambled of type $2$;
  4. (4)
    the distributionally proximal relation of $T$ is dense in $X\times X$ and $T$ is mean-L-unstable;
  5. (5)
    the distributionally proximal cell of $\mathbf{0}$ is dense in $X$ and there exists $x\in X$ and $\delta>0$ such that
    $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,\mathbf{0})>\delta\})>0.$
By Theorem 5.30 (5), we have the following consequence.
Corollary 5.31.
Let $T\in L(X)$. Assume that the distributionally proximal cell of $\mathbf{0}$ is dense. If $T$ has a non-trivial periodic point or an eigenvalue $\lambda$ with $|\lambda|\geq 1$, then $T$ admits a residual set of distributionally $m$-irregular points of type $2$ for some $m\in\mathbb{N}$.
Applying Theorem 5.30 we have the following characterization of distributional chaos of type $2$ for continuous linear operators on Fréchet spaces.
Theorem 5.32.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a distributionally semi-irregular point of type $2$;
  2. (2)
    $T$ admits a distributionally $m$-irregular points of type $2$ for some $m\in\mathbb{N}$;
  3. (3)
    $T$ is distributionally $m$-extremely $\delta$-scrambled of type $2$ for some $m\in\mathbb{N}$ and $\delta>0$;
  4. (4)
    the restriction of $T$ to some invariant closed $T$-invariant subspace $\widetilde{X}$ has a dense set of distributionally $m$-irregular vectors of type $2$ for some $m\in\mathbb{N}$.
Let $T\in L(X)$ and $m\in\mathbb{N}$. We say that $T$ is distributionally sensitive if there exists $\delta>0$ such that for any $x\in X$ and $\varepsilon>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon d(T^{n}x,T^{n}y)>\delta\})=1,$
and distributionally $m$-extremely sensitive if for any $x\in X$ and $\varepsilon,M>0$ there exists $y\in X$ with $d(x,y)<\varepsilon$ such that
$\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon p_{m}(T^{n}x-T^{n}y)>M\})=1.$
A vector $x\in X$ is called distributionally $m$-unbounded if there exists a subset $A$ of $\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}p_{m}(x)=\infty$.
The following result is essentially contained in [BBMP2013]*Propositions 7 and 8. One can also provide a proof using the idea in Propositions 5.24 and 5.29.
Proposition 5.33.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ is distributionally sensitive;
  2. (2)
    there exists $\delta>0$, a sequence $(y_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=\mathbf{0}$ and for any $k\in\mathbb{N}$
    $\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon d(T^{i}y_{k},\mathbf{0})>\delta\})\geq N_{k}(1-\tfrac{1}{k});$
  3. (3)
    there exists $m\in\mathbb{N}$, $\delta>0$, a sequence $(y_{k})_{k}$ in $X$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=\mathbf{0}$ and for any $k\in\mathbb{N}$
    $\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon p_{m}(T^{i}y_{k})>\delta\})\geq N_{k}(1-\tfrac{1}{k});$
  4. (4)
    $T$ is distributionally $m$-extremely sensitive for some $m\in\mathbb{N}$;
  5. (5)
    there exists $m\in\mathbb{N}$ and $y\in X$ with distributionally $m$-unbounded orbit;
  6. (6)
    there exists $m\in\mathbb{N}$ such that the set of all points $y\in X$ with distributionally $m$-unbounded orbits is residual in $X$.
According to Proposition 5.33, following the idea of the proof of Proposition 3.23, we have the following result.
Proposition 5.34.
Let $T\in L(X)$. If $T$ admits a distributionally chaotic pair of type $1$ then $T$ is distributionally sensitive.
Note that several sufficient conditions for dense distributional chaos for linear operators on separable Fréchet spaces are obtain in [BBMP2013]*Theorems 15 and 16. Combining Theorem 5.28 and Propositions 5.33 and 5.34, we have the following characterization of the existence of a dense set of distributionally semi-irregular points of type $1$ for continuous linear operators on Fréchet spaces.
Theorem 5.35.
Let $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributionally semi-irregular vectors of type $1$;
  2. (2)
    $T$ admits a residual set of distributionally $m$-irregular vectors for some $m\in\mathbb{N}$;
  3. (3)
    there exists $m\in\mathbb{N}$ such that for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally $m$-extremely scrambled of type $1$;
  4. (4)
    the distributionally proximal relation of $T$ is dense in $X\times X$ and $T$ is distributionally sensitive.
Applying Theorem 5.35, we have the following characterization of distributional chaos of type $1$ for continuous linear operators on Fréchet spaces. It should be noticed that the equivalences of (1)-(4) are already proved in [BBMP2013]*Theorem 12.
Theorem 5.36.
Let $T\in L(X)$. Then he following assertions are equivalent:
  1. (1)
    $T$ admits a distributionally semi-irregular vector of type $1$;
  2. (2)
    $T$ admits a distributionally chaotic pair of type $1$;
  3. (3)
    $T$ admits a distributionally irregular vector of type $1$;
  4. (4)
    $T$ is distributionally $m$-extremely chaotic of type $1$ for some $m\in\mathbb{N}$;
  5. (5)
    there restriction of $T$ to some invariant closed $T$-invariant subspace $X$ has a dense set of distributionally $m$-irregular vectors for some $m\in\mathbb{N}$.
A consequence of the Theorem 5.36 is the following result.
Proposition 5.37.
Let $T\in L(X)$. Then the set of distributionally $m$-irregular vectors for some $m\in\mathbb{N}$ is dense in the set of distributionally semi-irregular vectors of type $1$.
Following [BBMP2013], we say that $T\in L(X)$ satisfies the distributional chaos criterion if there exist sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ in $X$ such that
  1. (1)
    there exists $A\subset\mathbb{N}$ with $\operatorname{\overline{\mathrm{dens}}}(A)=1$ such that $\lim\limits_{A\ni n\to\infty}T^{n}x_{k}=\mathbf{0}$ for all $k\in\mathbb{N}$;
  2. (2)
    each $y_{k}$ is in $\overline{\operatorname{span}(\{x_{k}\colon k\in\mathbb{N}\})}$, $\lim_{k\to\infty}y_{k}=\mathbf{0}$ and there exists $\delta>0$ and an increasing sequence $(N_{k})_{k}$ in $\mathbb{N}$ such that for every $k\in\mathbb{N}$,
    $\operatorname{\mathrm{card}}(\{1\leq i\leq N_{k}\colon d(T^{i}y_{k},\mathbf{0})>\delta\})\geq N_{k}(1-\tfrac{1}{k}).$
Using the ideas developed in this subsection, we can give a new proof to the following result. Since the proof is similar to the one of Proposition 3.35, we leave it to the reader.
Proposition 5.38 ([BBMP2013]*Theorem 12).
Let $T\in L(X)$. Then $T$ is distributionally chaotic of type $1$ if and only if it satisfies the distributional chaos criterion.
For $i=1,2$, a vector subspace $Y$ of $X$ is called a distributionally $m$-irregular manifold of type $i$ for $T$ if every vector $y\in Y$ is distributionally $m$-irregular of type $i$ for $T$.
Similar to the proof of Theorem 3.36, we have the following two results. Note that a slight weak version of Proposition 5.40 was proved in [BBMP2013]*Theorem 15.
Proposition 5.39.
Let $X$ be a separable Fréchet space and $T\in L(X)$. If the distributionally asymptotic cell of $\mathbf{0}$ is dense in $X$, then either the distributionally asymptotic cell of $\mathbf{0}$ is $X$ or it has a dense distributionally $m$-irregular manifold of type $2$ for some $m\in\mathbb{N}$.
Proposition 5.40.
Let $X$ be a separable Fréchet space and $T\in L(X)$. If the distributionally asymptotic cell of $\mathbf{0}$ is dense in $X$, then $T$ is distributionally sensitive if and only if it has a dense distributionally $m$-irregular manifold of type $1$ for some $m\in\mathbb{N}$.
5.5. Distributional chaos for continuous linear operators on Banach spaces
The aim of this section is to slightly generalize the characterizations of distributional chaos for linear operators on Banach spaces in [BBMP2013] and [BBPW2018] via our new approach.
Combining [BBMP2013]*Proposition 8 and Propositions 5.15 and 5.33, we have the following important result.
Theorem 5.41.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then $T$ is mean-L-unstable if and only if it is distributionally sensitive.
Combining Theorems 5.30, 5.35 and 5.41, we have the following characterization about the existence of a dense set of distributionally irregular vectors for continuous linear operators on Banach spaces.
Theorem 5.42.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a dense set of distributionally semi-irregular vectors of type $2$;
  2. (2)
    $T$ admits a residual set of distributionally irregular vectors of type $1$;
  3. (3)
    for every sequence $(O_{j})_{j}$ of nonempty open subsets of $X$, there exists a sequence $(K_{j})_{j}$ of Cantor sets with $K_{j}\subset O_{j}$ such that $\bigcup_{j=1}^{\infty}K_{j}$ is distributionally extremely scrambled of type $1$;
  4. (4)
    the distributionally proximal relation of $T$ is dense in $X\times X$ and $T$ is mean-L-unstable;
  5. (5)
    the distributionally proximal relation of $T$ is dense in $X\times X$ and there exists $x\in X$ and $\delta>0$ such that
    $\operatorname{\overline{\mathrm{dens}}}(\{n\in\mathbb{N}\colon\lVert T^{n}x\rVert>\delta\})>0.$
By Theorem 5.42 (5), we have the following consequence.
Corollary 5.43.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Assume that the distributionally proximal cell of $\mathbf{0}$ is dense. If $T$ has a non-trivial periodic point or an eigenvalue $\lambda$ with $|\lambda|\geq 1$, then it is distributionally extremely chaotic of type $1$.
Applying Theorem 5.42, we have the following characterization of distributionally chaos for continuous linear operators on Banach spaces, which slightly strengths [BBMP2013]*Theorem 12 and [BBPW2018]*Theorem 2.
Theorem 5.44.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the following assertions are equivalent:
  1. (1)
    $T$ admits a distributionally semi-irregular vector of type $2$;
  2. (2)
    $T$ admits a distributionally chaotic pair of type $2$;
  3. (3)
    $T$ admits distributionally irregular vector of type $1$;
  4. (4)
    $T$ is distributionally extremely chaotic of type $1$;
  5. (5)
    the restriction of $T$ to some invariant closed $T$-invariant subspace $\widetilde{X}$ has a dense set of distributionally irregular vectors of type $1$.
A consequence of the Theorem 5.44 is the following result.
Proposition 5.45.
Let $(X,\lVert\cdot\rVert)$ be a Banach space and $T\in L(X)$. Then the set of distributionally irregular vectors of type $1$ is dense in the set of distributionally semi-irregular vectors of type $2$.
Combining Theorem 5.41, Propositions 5.39 and 5.40, we have the following result, which generalizes [BBPW2018]*Theorem 33.
Proposition 5.46.
Let $(X,\lVert\cdot\rVert)$ be a separable Banach space and $T\in L(X)$. If the distributionally asymptotic cell of $\mathbf{0}$ is dense in $X$, then either the distributionally asymptotic cell of $\mathbf{0}$ is $X$ or $T$ has a dense distributionally irregular manifold of type $1$.
Acknowledgments: J. Li (corresponding author) was supported in part by NFS of China (Grant Nos. 12222110 and 12171298).
{bibsection}
[…]